### v1

# AdaDec: Uncertainty-Guided Adaptive Decoding for LLM-based Code Generation

Abstract
Code generation with large language models (LLMs) is highly sensitive to token selection during decoding, particularly at uncertain decision points that influence program logic. While standard strategies like greedy and beam search treat all tokens uniformly, they overlook code-specific uncertainty patterns, leading to suboptimal performance. This paper presents an empirical study revealing that many generation errors stem from ranking mistakes at high-uncertainty steps, where the correct token is present but not top-ranked.
Motivated by these findings, we propose AdaDec, an uncertainty-guided adaptive decoding framework that integrates a token-level pause-then-rerank mechanism driven by token uncertainty (Shannon entropy). AdaDec learns model-specific uncertainty thresholds and applies a lookahead-based reranking strategy when uncertainty is high. Experiments on HumanEval and MBPP benchmarks show that AdaDec improves Pass1 accuracy by up to 15.5% over greedy decoding, outperforms or matches beam search, and reduces computational cost and latency through efficient, selective pausing. Our results highlight the promise of uncertainty-aware adaptive decoding for improving the reliability and efficiency of LLM-based code generation.
Index Terms:
Code Generation, Adaptive Strategy, Language Model
I Introduction
Large language models (LLMs) have significantly advanced the field of code generation by learning expressive representations of programming syntax and semantics from large-scale corpora of source code and documentation [1, 2, 3, 4, 5]. Models such as Code Llama [6] and DeepSeek-Coder [7]—pre-trained on billions of lines of code—exhibit strong capabilities in producing syntactically correct and context-aware code snippets from natural language prompts [8, 9, 10]. These capabilities have catalyzed a range of intelligent software engineering tasks, including code completion, automated bug fixing [11, 12, 13], and test-case generation, ultimately reducing developer effort and accelerating the software development process [14, 15].
At the core of these applications lies the decoding process, which translates learned representations into executable code sequences. Among various decoding strategies, greedy decoding and beam search are widely adopted due to their simplicity and efficiency. However, these methods often suffer from early commitment to suboptimal tokens, especially under uncertainty. Such premature decisions can severely degrade the quality of generated code, given the highly structured and error-intolerant nature of programming languages.
Decoding strategies, which determine token selection, are critical to the performance of LLMs in code generation. Greedy search and beam search, standard in natural language processing, are commonly applied to code generation but remain suboptimal due to code’s unique characteristics. Unlike natural language, code exhibits distinct token uncertainty patterns. Tokens such as identifiers or line-initial keywords are highly uncertain because their selection depends on future logic. In contrast, tokens constrained by syntax or semantics, for example the keyword in within loops or repeated variable references, are more deterministic. Furthermore, code correctness is highly sensitive to decisions made at these key uncertain points. An incorrect token can drastically change program behavior; for instance, choosing “for” versus “return” at the start of a line leads to divergent control flows. Despite the critical role of token uncertainty in code generation, greedy search ignores it completely, while beam search treats all decoding steps uniformly, overlooking crucial uncertain points and reducing generation efficiency.
We conduct an empirical study to better understand token-level uncertainty during the decoding process in LLM-based code generation. Using DeepSeek-Coder [7] and the HumanEval dataset [16], we analyze decoding failures by comparing model predictions that fail test cases with their corresponding ground-truth solutions. Finding-1: The analysis indicates that many errors arise from local ranking mistakes—where the correct token is present in the candidate set but not ranked first—resulting in logic drift. To investigate this phenomenon, we perform next-token prediction on 164 HumanEval problems, computing the Shannon entropy and the rank of the ground-truth token at each decoding step. This allows us to identify and characterize potential drift points during generation. Finding-2: The results suggest that entropy can serve as a useful signal for detecting logic drift and guiding adaptive pausing and reranking. However, due to variability across models, a fixed entropy threshold is suboptimal, highlighting the need for a dynamic, model-aware thresholding mechanism.
We introduce AdaDec, an uncertainty-guided adaptive decoding framework for code generation with large language models. AdaDec enhances standard decoding by incorporating a pause-then-rerank mechanism that reacts to model uncertainty, measured by Shannon entropy at each step. When uncertainty exceeds a threshold, decoding is paused and a reranking procedure is triggered. Because a fixed threshold does not generalize well across different models, AdaDec learns this threshold through a data-driven logistic regression approach. During reranking, AdaDec selects the top candidate tokens and evaluates them using a lookahead strategy inspired by A* search heuristics and a prior constrained generation method [17], scoring each candidate based on expected future continuations. The highest-scoring token is chosen for generation, and the process repeats until the sequence is complete.
We conduct extensive experiments on the HumanEval and MBPP benchmarks using models ranging from 0.6B to 8B parameters. Our uncertainty-guided adaptive decoding framework, AdaDec, consistently improves Pass1 accuracy by up to 15.5% over Greedy decoding and achieves results that are competitive with or better than Beam Search. In terms of efficiency, AdaDec selectively pauses decoding at a low rate (between 3.95% and 11.88%), which leads to reductions in token generation of up to 54.72% and sometimes decreases overall latency. Ablation studies confirm that the learned entropy threshold effectively identifies uncertain decoding steps, and both fixed and dynamic lookahead length selection strategies improve Pass1 over Greedy decoding across different models and datasets. Finally, AdaDec demonstrates strong generalizability, consistently improving code generation accuracy across a wide range of model sizes and both benchmark datasets.
In summary, this paper makes the following contributions:
  • An empirical study that reveals many code generation errors stem from local token ranking mistakes, and that Shannon entropy is a useful signal for identifying uncertain decoding steps where logic drift may occur.
  • An adaptive decoding approach named AdaDec, which uses an entropy-guided pause-then-rerank mechanism based on learned, model-specific thresholds and a lookahead strategy to improve generation quality.
  • Experimental results that show AdaDec improves Pass@1 accuracy by up to 15.5% over Greedy decoding, achieves high generation efficiency with reduced token count and latency, and generalizes effectively across diverse model sizes and benchmark datasets.
  • All code and are publicly available at [18].
II Empirical Study
We conduct an empirical study to investigate the characteristics of token uncertainty in LLM-based code generation and analyze the feasibility of integrating token uncertainty into decoding process. Specifically, the study aims to answer the following research questions:
  • RQ1: How does token uncertainty affect the behavior of generated code?
  • RQ2: How can token uncertainty be leveraged to improve token selection during decoding?
II-A Study Setup
II-A1 Models
To ensure the generality of our findings, we evaluate AdaDec on a diverse set of open-source code generation models with varying scales and architectures. Specifically, we include:
  • DeepSeek-Coder-1.3B-Instruct (DS-1.3B), DeepSeek-Coder-6.7B-Instruct (DS-6.7B) [7]: These models are trained from scratch on 2 trillion tokens comprising 87% code and 13% natural language, employing a fill-in-the-blank objective and a 16K context window to support project-level code completion and infilling tasks.
  • StableCode-Instruct-3B (ST-3B) [19]: A 2.7B parameter decoder-only model fine-tuned using Direct Preference Optimization (DPO), demonstrating state-of-the-art performance among similarly sized models on MultiPL-E and MT-Bench coding benchmarks.
  • CodeLlama-7B-Python-hf (CL-7B) [6]: A 7B parameter model specialized for Python code generation, offering capabilities such as code completion, infilling, and instruction following, and achieving strong performance on HumanEval and MBPP benchmarks.
  • Qwen3-0.6B (QW-0.6B), Qwen3-1.7B (QW-1.7B), Qwen3-4B (QW-4B), Qwen3-8B (QW-8B) [20]: A family of multilingual large language models ranging from 0.6B to 8B parameters, featuring dynamic mode switching between reasoning and fast-response modes, and supporting a wide range of tasks including code generation.
These models span from 0.6B to 8B parameters and are accessed via HuggingFace Transformers. They cover a range of architectures and instruction-tuning schemes, ensuring a diverse evaluation across different model families and sizes.
II-A2 Implementation
Our implementation is based on PyTorch and HuggingFace Transformers, and decoding is performed in a batch-free, single-example fashion to avoid interference between instances. All models are run using half-precision (i.e., torch.float16) to accelerate inference and reduce memory usage. The maximum number of generated tokens is capped at 512 for all methods to ensure consistent comparison across models.
Experiments are conducted on a machine equipped with an NVIDIA RTX 4090D GPU (24GB VRAM), an Intel(R) Xeon(R) Platinum 8474C CPU with 16 vCPUs, and 80GB of system memory.
II-B RQ1: How does token uncertainty affect the behavior of generated code?
To address this research question, we first use DS-6.7B as a representative model to generate code for 164 programming problems in the widely used HumanEval dataset [16]. We then investigate the phenomenon of logic drifts during decoding, which often lead to incorrect behavior in the generated code. Finally, we explore the relationship between logic drifts and token-level uncertainty, quantified using Shannon entropy, a commonly used metric for estimating uncertainty [21].
II-B1 Method
For each programming problem in the HumanEval dataset, we use DS-6.7B to generate code snippets using a greedy decoding strategy. We then filter out the generations that successfully pass all test cases. Based on an initial human analysis, we observe that in several cases, the model begins with a correct sequence but introduces a critical error—often in the form of incorrect code structures—at a specific step. At these error points, the correct next token is not missing from the candidate set but is ranked below the top-1 position (commonly second or third), resulting in an irreversible deviation from the correct logic. This suggests that some generation failures are caused by local ranking errors at specific decoding steps, ultimately leading to incorrect code behavior. We refer to such steps as logic drift points during decoding and aim to identify them.
Given the high cost of fully manual annotation, we develop a semi-automatic approach that combines abstract syntax tree (AST) comparison with human verification and fine-grained token-level annotation. To ensure annotation focuses only on faulty generations, we first filter out all generated samples that can pass the tests. Subsequently, we construct ASTs for both the ground-truth and generated code, and normalize them by removing identifiers and comments. This normalization ensures that the comparison focuses solely on structural differences. We then perform a parallel traversal to compare the two normalized ASTs node by node. A mismatch is flagged when node types differ, operators or constants are inconsistent, or discrepancies occur in control structures. The mismatch is then traced back to its corresponding line in original code and extracted as a candidate drift point. This candidate is then reviewed by a human annotator to confirm whether it represents an actual drift point and to annotate the precise token where the drift begins. For example, a mismatch between the generated line if shift > len(s): and the ground-truth line if shift >= len(digits): triggers a drift point at the position of the operator >=.
In addition to the identified drift points, we collect all other non-drift decoding steps from the same generated code snippets. We compute the Shannon entropy for each step and compare the entropy distribution of drift points with that of non-drift steps. This comparison allows us to investigate the relationship between logic drift and token uncertainty, as measured by Shannon entropy. Given a probability distribution $\bm{p}$, the corresponding entropy $H(\bm{p})$ is calculated as follows:
$H(\bm{p})=-\sum_{p_{i}\in\bm{p}}p_{i}\log p_{i}$
II-B2 Results
We identify a total of 46 drift points and 2,889 non-drift decoding steps across 46 HumanEval problems. Figure 1 presents a boxplot comparing the entropy distribution at drift points with that of all non-drift steps. The entropy values at drift points are markedly higher: the median entropy is 1.26, with a mean of 1.39, compared to a much lower median of 0.03 and mean of 0.30 for the rest. The interquartile range (IQR) for drift points is also shifted higher (0.63–2.01) relative to that of non-drift steps (0.003–0.34). The results show that tokens at drift points exhibit significantly higher entropy values compared to the rest, indicating elevated uncertainty at these critical steps.
Figure 1: Entropy comparison between drift points and non-drift decoding steps.
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Finding 1: This empirical pattern suggests that token uncertainty, as measured by Shannon entropy, may serve as a useful signal for identifying potential logic drifts and predicting code behavior deviations during the decoding process.
II-C RQ2: How can token uncertainty be leveraged to improve token selection during decoding?
To integrate token uncertainty into the token selection process and enhance the likelihood of generating correct code, we further investigate the relationship between entropy and the rank positions of expected tokens during decoding.
II-C1 Method
We perform next-token prediction using 164 problems and their corresponding ground-truth solutions from the HumanEval dataset. Specifically, given a problem $P$ and its ground-truth solution $S$, we tokenize them into sequences $\mathrm{T}^{P}$ and $\mathrm{T}^{S}$, respectively. At each decoding step $k$, we feed the concatenated sequence $\mathrm{T}^{P}\oplus\mathrm{T}^{S}_{:k}$ into the LLM to obtain a probability distribution over the next token. Here, $\mathrm{T}^{S}_{:k}$ denotes the prefix of the ground-truth solution containing the first $k$ tokens, and $\oplus$ indicates sequence concatenation. By iterating over all positions $k$ from 1 to the length of $\mathrm{T}^{S}$, we collect a set of predicted probability distributions. For each prediction, we compute the Shannon entropy of the distribution and record the rank of the expected (ground-truth) token within it.
We apply this next-token prediction procedure to a range of open-source LLMs, including four code-specific models, DS-1.3B, DS-6.7B, ST-3B, and CL-7B, as well as four general-purpose models from the Qwen3 family: QW-0.6B, QW-1.7B, QW-4B, and QW-8B. For Qwen3, we disable the reasoning-specific settings (e.g., thinking mode). In total, we collect 180,879 token-level probability distributions for two DeepSeek-Coder models, 169,496 for StableCode, 177,650 for CodeLlama, and 129,165 for Qwen3 series.
II-C2 Results
We compute the Spearman correlation between entropy and the rank of the ground-truth token across all models. As shown in Table I, the results consistently exhibit a positive correlation, further reinforcing the notion that entropy serves as a meaningful signal for identifying uncertain decoding steps. We omit p-values from the table, as the large sample sizes (approximately 130,000 to 180,000 tokens per model) yield statistically significant results with p-values effectively indistinguishable from zero.
Table I: Spearman correlation $\bm{\rho}$ between entropy and the rank of ground-truth tokens.
Model DS-1.3B DS-6.7B ST-3B CL-7B QW-0.6B QW-1.7B QW-4B QW-8B
$\bm{\rho}$ 0.4728 0.4595 0.4620 0.4512 0.5249 0.5028 0.4936 0.4877
Figure 2 shows how the percentage of decoding steps with entropy exceeding a given threshold changes as the threshold increases. Figure 3 illustrates the corresponding change in the average rank of ground-truth tokens, separated by whether the entropy at each step is above or below the threshold. These plots reveal that different models exhibit varying sensitivity to entropy thresholds. For example, at a threshold of 1, CL-7B and QW-0.6B have approximately 15% of decoding steps exceeding the threshold, while other models remain below 10%. In terms of average rank, two models (DS-6.7B and QW-1.7B) exceed a value of 4.5, one model (CL-7B) is around 2.6, and the rest fall in the range of 3.5 to 4.
Figure 2: Change in the percentage of decoding steps exceeding a given entropy threshold as the threshold increases.
Figure 3: Change in the average rank of ground-truth tokens above and below a given entropy threshold, as the threshold increases.
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Finding 2: The observed correlation between entropy and the rank of the ground-truth token suggests that entropy can be used as an indicator to adaptively pause the decoding process and rerank uncertain tokens. However, our entropy percentile analysis shows that it is difficult to define a universal, fixed entropy threshold across all models that effectively balances pause frequency and the number of reranking candidates. This highlights the need for a dynamic mechanism to determine the model-specific entropy threshold.
III Approach
We now introduce AdaDec, an uncertainty-guided adaptive decoding framework for LLM-based code generation. AdaDec integrates a token-level pause-then-rerank mechanism into the decoding process of the target LLM, drawing on insights from our empirical analysis.
III-A Overview
Figure 4: Approach Overview of AdaDec
Given a target model LM, the adaptive decoding process of AdaDec, which incorporates the pause-and-rerank mechanism, is illustrated in Figure 4. At each decoding step $t$, after LM predicts a probability distribution $\bm{p}_{t}$ for the next token, AdaDec determines whether to pause the decoding by checking if the entropy $H(\bm{p}_{t})$ of $\bm{p}_{t}$ exceeds a model-specific threshold $\tau^{\bm{LM}}$. If $H(\bm{p}_{t})$ surpasses $\tau^{\bm{LM}}$, indicating high uncertainty, AdaDec triggers the reranking process. Specifically, it selects the top-$K$ candidate tokens from the distribution and applies a lookahead strategy, re-scoring each token’s probability based on its expected likelihood over the next $N$ decoding steps. The token with the highest re-scored probability is then selected as the next token. This process continues iteratively until an end-of-sequence (EOS) token is generated.
The details of the pause checking and lookahead reranking are described in the following sections.
III-B Entropy-based Pause Checking
We use the Shannon entropy $H(\bm{p}_{t})$ over the probability distribution $\bm{p}_{t}$ predicted by the model LM to measure the token uncertainty. When $H$ exceeds a threshold, AdaDec pauses the decoding process. As revealed in the empirical study, a fixed entropy threshold may not generalize well across models with different output distributions. Therefore, we propose a data-driven method to learn a model-specific threshold $\tau^{\bm{LM}}$ using logistic regression.
III-B1 Data Collection
To enable threshold learning, we cast the problem as a binary classification task. Each sample corresponds to a decoding step and is labeled based on whether the ground truth token is ranked top-1 in the predicted distribution. The input feature for classification is the entropy $H(\bm{p}_{t})$, which reflects the model’s uncertainty at that step.
We generate training data by conducting ground-truth–guided next-token prediction with the target model LM on the BigCodeBench dataset [22]. In this process, the ground truth token is provided as input at each step instead of the model’s own prediction. During this process, we record the model’s predicted probability distribution over the vocabulary at each decoding step and calculate the entropy of the predicted distribution. Based on whether the ground truth token is ranked as the top-1 prediction, we assign a binary label: 1 if it is top-1, and 0 otherwise. Based on this process, we collect between approximately 100,000 and 160,000 positive samples (labeled as 1) and between 17,000 and 23,000 negative samples (labeled as 0) per model. For example, in the case of DS-6.7B, we obtain 161,856 positive samples and 19,023 negative samples. The substantial imbalance between the number of positive and negative samples further highlights the inefficiency of beam search, which performs reranking at every decoding step regardless of model uncertainty.
We then balance the dataset by randomly downsampling the positive samples and selecting an equal number of negative samples, which is a common technique for handling imbalanced classification tasks [23]. This choice is driven by the nature of logic drift in code generation: once a drift point occurs, it typically leads to irreversible semantic errors. Therefore, it is essential to maximize the classifier’s ability to detect negative cases, even if they are relatively scarce. By constructing a balanced dataset, we prevent the model from being biased toward the majority class and enable it to more effectively learn the decision boundary for identifying uncertain decoding steps. This strategy enhances the classifier’s sensitivity to potential drift points, thereby improving the overall reliability of the entropy-based triggering mechanism.
III-B2 Regression Model Training
We train a logistic regression model to predict whether the ground truth token is the top-1 prediction, using only the entropy value as input. The model estimates the probability of a positive label using the sigmoid function:
$P(y=1\mid H^{(i)})=\frac{1}{1+\exp(-(\beta_{0}+\beta_{1}H^{(i)}))},$
where $H^{(i)}$ denotes the entropy of the $i$-th sample, and $\beta_{0}$, $\beta_{1}$ are the model parameters to be learned. These parameters are learned by minimizing the binary cross-entropy loss through maximum likelihood estimation. Note that, since the training data is collected using the target LM, the trained logistic regression model is specific to that LM.
III-B3 Threshold Estimation
To estimate a model-specific entropy threshold $\tau^{\bm{LM}}$, we use the trained logistic regression model to perform binary classification on the validation set, where the task is to predict whether the top-1 token matches the ground-truth token. As the logistic regression model outputs probabilities, we first identify the optimal probability threshold that maximizes classification accuracy, and then convert this probability threshold into an equivalent entropy threshold.
Specifically, we conduct a grid search over probability thresholds in the range [0.01, 0.99], using a step size of 0.01. For each candidate threshold, a sample is classified as positive if its predicted probability exceeds the threshold, and negative otherwise. The threshold that achieves the highest classification accuracy on the validation set is selected as the optimal probability threshold, denoted by $p^{*}$. We then convert this optimal probability threshold into the corresponding entropy threshold using the parameters of the trained logistic regression model. Given the intercept $\beta_{0}$ and the coefficient $\beta_{1}$, the entropy threshold $\tau^{\bm{LM}}$ is computed using the log-odds transformation:
$\tau^{\bm{LM}}=\frac{\log\left(\frac{p^{*}}{1-p^{*}}\right)-\beta_{0}}{\beta_{1}}.$
Entropy values below $\tau^{\bm{LM}}$ are classified as positive (i.e., the model is likely to correctly predict the top-1 token), while those above are classified as negative. This entropy threshold is then used as the final decision boundary for the target language model during decoding.
III-B4 Pause Triggering
During decoding, AdaDec uses the learned threshold $\tau^{\bm{LM}}$ to decide when to trigger candidate reranking. If the entropy at a decoding step exceeds $\tau^{\bm{LM}}$, it indicates high uncertainty, suggesting that the top-1 prediction may be incorrect. In such cases, the decoding process is paused, and the top candidates are reranked to improve accuracy.
III-C Lookahead-based Token Reranking
AdaDec employs a lookahead re-scoring strategy for token reranking, enhancing selection by assessing the future impact of each candidate token at every decoding step. Inspired by A* search heuristics and prior work on constrained generation [17], this method extends token probability evaluation beyond immediate next-token predictions to consider longer-term sequence quality. The process is governed by two hyperparameters: the lookahead beam size ($B$), which determines the number of candidate tokens considered, and the lookahead length ($L$), which sets the length of the future token sequence evaluated.
III-C1 Lookahead Process
At decoding step $t$, the model outputs a probability distribution $\bm{p}_{t}$ over the vocabulary for the next token. From this distribution, we select the top-$B$ tokens to form the candidate set $\mathcal{C}_{t}=\{c_{t}^{(1)},c_{t}^{(2)},\dots,c_{t}^{(B)}\}$.
For each candidate token $c_{t}^{(k)}$, we construct a lookahead trajectory of length $L$ by appending $L$ tokens generated via greedy decoding, conditioned on the current input and previously generated tokens. This results in the trajectory $\mathbf{T}_{t}^{(k)}$:
$\mathbf{T}_{t}^{(k)}=(c_{t}^{(k)},\hat{c}_{t+1}^{(k)},\hat{c}_{t+2}^{(k)},\dots,\hat{c}_{t+L}^{(k)}),$
where each $\hat{c}_{t+i}^{(k)}$ is the most probable token predicted at step $t+i$, given the prefix ending with $c_{t}^{(k)}$.
Trajectory Scoring. We evaluate each trajectory by computing the average log-probability of its tokens:
$\text{score}(c_{t}^{(k)})=\frac{1}{L+1}\left[\log P(c_{t}^{(k)})+\sum_{i=1}^{L}\log P(\hat{c}_{t+i}^{(k)})\right],$
where $P(\cdot)$ denotes the model’s predicted probability for a token. To ensure a fair comparison across candidates with varying lookahead lengths (as discussed in the next section on lookahead length selection), we normalize the total log-probability by computing the geometric mean of token probabilities in the above equation, rather than relying solely on log-likelihood. This normalization mitigates the bias against longer sequences, which tend to accumulate lower joint probabilities. The resulting score captures the model’s overall confidence in the full trajectory, balancing immediate token likelihood with the coherence of future predictions.
Token Selection. The candidate with the highest trajectory score is selected as the output token at step $t$:
$c^{*}_{t}=\arg\max_{c_{t}^{(k)}\in\mathcal{C}_{t}}\text{score}(c_{t}^{(k)}).$
This process is repeated at each decoding step, ensuring that token selection is guided not just by immediate probabilities, but by the expected quality of the downstream sequence.
This lookahead mechanism improves generation quality by favoring tokens that lead to fluent and coherent continuations, which is particularly important in structured domains like code generation. The hyperparameters $B$ and $L$ control the trade-off between computational cost and generation quality: larger values offer better foresight at the expense of increased inference time.
III-C2 Lookahead Beam Size Selection
To determine an appropriate value for $B$, we analyze the training data collected in Section III-B1. Specifically, we examine the average rank of the ground truth token at decoding steps where the Shannon entropy is either above or below the threshold $\tau^{\bm{LM}}$, which is learned via logistic regression. After filtering out samples where the ground truth token’s rank exceeds 20, we observe a clear distinction: when the entropy is below or equal to the threshold, the average rank typically falls between 1.1 and 1.3, indicating high model confidence. In contrast, when the entropy exceeds the threshold—thus triggering a pause—the average rank increases to between 2.6 and 3.4.
These findings support the use of a small beam size (e.g., $B=3$) to achieve a practical trade-off between candidate coverage and computational efficiency.
III-C3 Lookahead Length Selection
We explore two strategies for determining the lookahead length $L$ during trajectory scoring:
  • Fixed: A simple approach that uses a constant lookahead length (e.g., 5) for all candidate tokens.
[…]
The line-aware dynamic strategy allows for adaptive lookahead tailored to the syntactic structure of code, potentially improving alignment with natural coding boundaries. In Section IV, we empirically compare the two strategies.
IV Evaluation
We conduct extensive experiments to evaluate the effectiveness of AdaDec. Specifically, we aim to answer the following research questions:
  • RQ1 (Effectiveness): How effective is AdaDec in improving the functional correctness of generated code?
  • RQ2 (Efficiency): How efficient is AdaDec in terms of computational cost?
  • RQ3 (Ablation Study): How effective is the logistic regression–based threshold learning component, and how do the two lookahead length selection strategies affect performance?
  • RQ4 (Generalizability): How well does AdaDec generalize across different models and datasets?
IV-A Experimental Setup
Datasets. We evaluate our method on two widely used code generation benchmarks. The first is HumanEval [16], which contains 164 hand-crafted problems. The second is MBPP [26], from which we use the first 200 problems for evaluation.
Baselines. We compare AdaDec against two standard decoding strategies: Greedy decoding and Beam Search with a beam size of 3 (equal to the lookahead beam size used in AdaDec). These baselines represent common trade-offs between generation quality and computational efficiency.
Variants of AdaDec. To examine the impact of different lookahead length (L) strategies, we evaluate two variants of our method:
  • $\textsc{AdaDec}_{L=5}$: Applies the fixed lookahead length of $L=5$ for all reranking steps.
  • $\textsc{AdaDec}_{dynL}$: Uses the dynamic, line-aware lookahead length with a limit of 10.
IV-B RQ1: Effectiveness in Improving Functional Correctness
Table II presents the Pass@1 performance of four decoding strategies, Greedy, Beam Search, $\textsc{AdaDec}_{L=5}$, and $\textsc{AdaDec}_{dynL}$, across eight LLMs evaluated on the HumanEval and MBPP datasets.
Table II: Pass@1 comparison of Greedy, Beam Search, $\textsc{AdaDec}_{L=5}$, and $\textsc{AdaDec}_{dynL}$ on HumanEval and MBPP. Values in parentheses indicate the change relative to Greedy. The “AVG. $\Delta$” row reports the average difference from Greedy across all eight models.
Model HumanEval MBPP
Greedy Beam Search $\textsc{AdaDec}_{L=5}$ $\textsc{AdaDec}_{dynL}$ Greedy Beam Search $\textsc{AdaDec}_{L=5}$ $\textsc{AdaDec}_{dynL}$
DS-1.3B 56.71% 59.76% (+3.05%) 64.02% (+7.32%) 61.59% (+4.88%) 36.50% 40.50% (+4.00%) 37.00% (+0.50%) 37.00% (+0.50%)
DS-6.7B 71.95% 74.39% (+2.44%) 74.39% (+2.44%) 70.73% (-1.22%) 47.50% 51.00% (+3.50%) 51.00% (+3.50%) 52.50% (+5.00%)
ST-3B 55.49% 55.49% (+0.00%) 55.49% (+0.00%) 55.49% (+0.00%) 35.00% 37.50% (+2.50%) 40.00% (+5.00%) 40.00% (+5.00%)
CL-7B 38.41% 42.68% (+4.27%) 40.24% (+1.83%) 42.68% (+4.27%) 35.00% 36.50% (+1.50%) 38.50% (+3.50%) 40.00% (+5.00%)
QW-0.6B 21.34% 28.05% (+6.71%) 25.00% (+3.66%) 25.61% (+4.27%) 12.00% 20.50% (+8.50%) 18.00% (+6.00%) 19.00% (+7.00%)
QW-1.7B 46.34% 46.34% (+0.00%) 50.00% (+3.66%) 47.56% (+1.22%) 34.00% 40.00% (+6.00%) 37.00% (+3.00%) 37.00% (+3.00%)
QW-4B 57.32% 65.85% (+8.54%) 64.02% (+6.71%) 67.07% (+9.76%) 23.50% 35.00% (+11.50%) 39.00% (+15.50%) 37.00% (+13.50%)
QW-8B 61.59% 66.46% (+4.88%) 71.34% (+9.76%) 70.12% (+8.54%) 40.50% 49.00% (+8.50%) 49.00% (+8.50%) 50.50% (+10.00%)
AVG. $\Delta$ +3.73% +4.42% +3.96% +5.75% +5.69% +6.12%
Comparison with Greedy Search. Both variants of our adaptive decoding approach, $\textsc{AdaDec}_{L=5}$ and $\textsc{AdaDec}_{dynL}$, consistently outperform Greedy decoding across nearly all models and datasets, with the exception of $\textsc{AdaDec}_{dynL}$ on DS-6.7B for HumanEval. On HumanEval, $\textsc{AdaDec}_{L=5}$ yields an average absolute improvement of +4.42%, while $\textsc{AdaDec}_{dynL}$ follows closely with +3.96%, compared to the Greedy baseline. The performance gains are even more prominent on MBPP, where $\textsc{AdaDec}_{dynL}$ achieves the highest average improvement of +6.12%, with $\textsc{AdaDec}_{L=5}$ closely behind at +5.69%.
Notably, certain models show substantial improvements. For instance, QW-8B exhibits a +9.76% increase with $\textsc{AdaDec}_{L=5}$ on HumanEval and a +10.00% increase with $\textsc{AdaDec}_{dynL}$ on MBPP. Similarly, QW-4B sees large gains across both datasets (ranging from +6.71% to +15.50%), underscoring the effectiveness of adaptive decoding especially for mid- to large-sized models.
Comparison with Beam Search. When compared to Beam Search, the performance of our adaptive decoding strategies is competitive and often superior, particularly on MBPP. On HumanEval, $\textsc{AdaDec}_{L=5}$ slightly outperforms Beam Search on average (+4.42% vs. +3.73% over Greedy), while $\textsc{AdaDec}_{dynL}$ achieves comparable gains (+3.96%). On MBPP, $\textsc{AdaDec}_{dynL}$ demonstrates the best average improvement over Greedy (+6.12%), surpassing both Beam Search (+5.75%) and $\textsc{AdaDec}_{L=5}$ (+5.69%).
While Beam Search excels in some individual cases (e.g., DS-6.7B and CL-7B on HumanEval), adaptive decoding methods show more robust performance across both datasets and models. $\textsc{AdaDec}_{L=5}$ and $\textsc{AdaDec}_{dynL}$ each achieve the highest Pass@1 scores in 6 out of the 16 evaluation settings (2 datasets × 8 models), highlighting their consistent generalization advantage across tasks and model sizes. Moreover, our adaptive decoding approach offers improved efficiency, as discussed in the next section (RQ2).
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Summary: Our uncertainty-guided adaptive decoding approach, AdaDec, outperforms Greedy decoding on average by up to +6.12% in Pass@1 across HumanEval and MBPP, and achieves competitive or superior results compared to Beam Search. $\textsc{AdaDec}_{dynL}$ delivers the highest gains on MBPP, while $\textsc{AdaDec}_{L=5}$ leads on HumanEval.
IV-C RQ2: Efficiency in Computational Cost
We evaluate the efficiency of AdaDec in terms of computational cost on the HumanEval and MBPP datasets from the following perspectives. First, we measure the pause rate of AdaDec to assess the possible overhead introduced by its pause-then-rank mechanism. Second, we compare the count of generated tokens and the prediction latency of AdaDec against Greedy decoding and Beam Search.
Pause Rate. As presented in Table III, the pause rate ranges from 3.95% to 11.88%, with an average of 7.15%. This low activation frequency indicates that AdaDec concentrates computational effort on a small subset of high-uncertainty steps, effectively balancing local correction with overall efficiency. In all experiments, the pause rate remains below 12%, demonstrating that selective triggering successfully avoids the high cost associated with global lookahead.
Table III: Pause ratio of the models on HumanEval and MBPP.
DS-1.3B DS-6.7B ST-3B CL-7B QW-0.6B QW-1.7B QW-4B QW-8B
$\textsc{AdaDec}_{L=5}$
HumanEval 6.54% 5.53% 4.76% 3.95% 6.10% 11.07% 7.95% 5.98%
MBPP 8.32% 5.82% 6.35% 5.20% 6.33% 11.41% 10.64% 8.41%
$\textsc{AdaDec}_{dynL}$
HumanEval 6.69% 5.60% 4.72% 3.95% 5.18% 10.06% 9.37% 6.39%
MBPP 8.28% 5.89% 6.51% 5.45% 6.21% 11.88% 10.18% 7.96%
Token Count. Table IV summarizes the results. A consistent and noteworthy outcome across both datasets is that AdaDec substantially reduces the average number of generated tokens. This reduction is not achieved by truncating outputs or constraining sequence length, but rather through more focused and efficient content selection during decoding. By leveraging its pause-and-rank mechanism to identify promising continuations early, AdaDec avoids veering into verbose sequences.
Compared with Greedy decoding, representative examples include: on MBPP, QW-0.6B reduces token count by 54.72%, from 341.07 to 154.45 with $\textsc{AdaDec}_{dynL}$; on HumanEval, CL-7B drops from 104.92 to 88.43 tokens (-15.71%) with $\textsc{AdaDec}_{dynL}$; and on MBPP, both ST-3B and QW-1.7B achieve 24–34% reduction swith $\textsc{AdaDec}_{L=5}$. These results highlight AdaDec’s built-in “generation compression” effect: by pausing to rank and select more meaningful continuations, the model often generates correct code using fewer tokens.
Latency. As shown in Table V, AdaDec sometimes reduces total generation latency, despite the additional overhead of pauses and candidate reranking. This counterintuitive outcome stems from the shorter outputs produced, which in many cases offset, or even outweigh, the cost of reranking. On MBPP, for instance, $\textsc{AdaDec}_{L=5}$ reduces the average latency for Qwen3-0.6B by 31.12% (from 6.92 s to 4.76 s), for Stable-3B by 15.53% (1.17 s to 0.99 s), and for CodeLlama-7B by 18.10% (3.23 s to 2.65 s). On HumanEval, $\textsc{AdaDec}_{dynL}$ achieves a 13.45% reduction for Qwen3-0.6B (3.67 s to 3.17 s).
Regarding Beam Search, although its latency is comparable to AdaDec, its computational cost is substantially higher. This is because Beam Search performs many more logits computations and token predictions during decoding. While these computations are parallelized on CPUs or GPUs and thus do not significantly increase latency, the added computational overhead results in greater energy consumption.
Table IV: Comparison of the average number of tokens generated per problem by Greedy, Beam Search, $\textsc{AdaDec}_{L=5}$, and $\textsc{AdaDec}_{dynL}$ on HumanEval and MBPP (the brackets indicate the change ratio relative to Greedy)
Model HumanEval MBPP
Greedy Beam Search $\textsc{AdaDec}_{L=5}$ $\textsc{AdaDec}_{dynL}$ Greedy Beam Search $\textsc{AdaDec}_{L=5}$ $\textsc{AdaDec}_{dynL}$
DS-1.3B 66.27 74.80 (+12.86%) 64.38 (-2.86%) 66.59 (+0.47%) 75.61 71.10 (-5.96%) 54.49 (-27.93%) 59.36 (-21.49%)
DS-6.7B 64.32 67.68 (+5.21%) 62.10 (-3.45%) 60.76 (-5.55%) 70.25 62.18 (-11.48%) 59.28 (-15.62%) 53.54 (-23.78%)
ST-3B 79.98 73.23 (-8.44%) 70.18 (-12.25%) 68.10 (-14.85%) 71.83 63.12 (-12.13%) 54.66 (-23.91%) 53.53 (-25.48%)
CL-7B 104.92 104.65 (-0.26%) 90.04 (-14.19%) 88.43 (-15.71%) 90.88 81.03 (-10.83%) 68.21 (-24.94%) 68.61 (-24.50%)
QW-0.6B 168.62 98.71 (-41.46%) 107.54 (-36.22%) 100.10 (-40.64%) 341.07 80.94 (-76.27%) 171.21 (-49.80%) 154.45 (-54.72%)
QW-1.7B 100.57 73.57 (-26.84%) 81.89 (-18.57%) 83.40 (-17.07%) 131.80 85.45 (-35.17%) 86.38 (-34.46%) 98.66 (-25.15%)
QW-4B 115.34 89.61 (-22.31%) 111.87 (-3.01%) 89.08 (-22.77%) 86.55 52.50 (-39.35%) 82.72 (-4.43%) 74.22 (-14.25%)
QW-8B 95.24 81.85 (-14.06%) 84.79 (-10.97%) 89.10 (-6.45%) 94.58 60.25 (-36.30%) 78.34 (-17.18%) 71.96 (-23.91%)
Table V: Comparison of the average generation time per problem (in seconds) by Greedy, Beam Search, $\textsc{AdaDec}_{L=5}$, and $\textsc{AdaDec}_{dynL}$ on HumanEval and MBPP (the brackets indicate the change ratio relative to Greedy)
Model HumanEval MBPP
Greedy Beam Search $\textsc{AdaDec}_{L=5}$ $\textsc{AdaDec}_{dynL}$ Greedy Beam Search $\textsc{AdaDec}_{L=5}$ $\textsc{AdaDec}_{dynL}$
DS-1.3B 1.18 2.08 (+76.79%) 1.51 (+28.33%) 1.84 (+56.35%) 1.14 1.43 (+25.97%) 1.15 (+1.54%) 1.77 (+55.76%)
DS-6.7B 2.47 6.27 (+154.15%) 3.54 (+43.53%) 4.10 (+66.40%) 2.12 2.97 (+40.03%) 2.28 (+7.62%) 2.64 (+24.38%)
ST-3B 1.50 2.87 (+91.14%) 1.95 (+29.73%) 2.18 (+45.27%) 1.17 1.48 (+26.57%) 0.99 (-15.53%) 1.35 (+15.83%)
CL-7B 5.02 13.69 (+172.58%) 5.72 (+13.85%) 6.41 (+27.75%) 3.23 5.28 (+63.21%) 2.65 (-18.10%) 3.65 (+12.93%)
QW-0.6B 3.67 3.04 (-16.93%) 3.50 (-4.41%) 3.17 (-13.45%) 6.92 2.07 (-70.10%) 4.76 (-31.12%) 5.72 (-17.35%)
QW-1.7B 2.70 2.71 (+0.60%) 3.34 (+23.71%) 4.05 (+50.31%) 2.83 2.57 (-9.30%) 2.84 (+0.16%) 5.46 (+92.86%)
QW-4B 4.19 7.20 (+72.05%) 7.32 (+74.89%) 7.44 (+77.66%) 2.71 2.10 (-22.62%) 4.03 (+48.81%) 5.15 (+89.99%)
QW-8B 5.05 9.37 (+85.51%) 7.01 (+38.77%) 10.12 (+100.46%) 4.03 3.44 (-14.74%) 4.85 (+20.23%) 5.69 (+41.00%)
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Summary: AdaDec selectively pauses decoding at a low rate (3.95–11.88%), reducing token count by up to 54.72% and frequently lowering latency, thereby demonstrating both efficient and accurate decoding.
IV-D RQ3: Ablation Study
We perform an ablation study to evaluate the key components of AdaDec. Specifically, we first investigate the effectiveness of the learned entropy threshold $\tau^{\bm{LM}}$ in identifying top-1 ground-truth tokens based on entropy. Then, we compare two variants of AdaDec to analyze the impact of different lookahead length selection strategies.
Effectiveness of Learned Entropy Thresholds. Table VI summarizes the performance of a logistic regression classifier trained with learned entropy thresholds ($\tau^{LM}$) across various language models. Overall, the classifier demonstrates strong effectiveness in predicting whether the top-1 token matches the ground truth. The highest performance is observed on DS-6.7B, achieving an accuracy of 91.05%, an F1 score of 0.9512, and an AUC of 0.9305. Other models, including DS-1.3B, ST-3B, and CL-7B, also perform well, with F1 scores above 0.94 and AUC values exceeding 0.89. Notably, CL-7B exhibits the highest learned entropy threshold ($\tau^{LM}=1.9353$), suggesting a greater tolerance. These results support the effectiveness of learned entropy thresholds as a reliable and lightweight signal for triggering pauses and reranking.
Table VI: Evaluation of logistic regression models trained to predict whether the ground truth token is the top-1 prediction. The learned $\bm{\tau}^{\bm{LM}}$ denotes the entropy threshold selected via logistic regression.
Model Accuracy Precision Recall F1 Score AUC Learned $\bm{\tau}^{\bm{LM}}$
DS-1.3B 0.8994 0.9251 0.9640 0.9442 0.9268 0.9850
DS-6.7B 0.9105 0.9276 0.9761 0.9512 0.9305 1.1338
ST-3B 0.9038 0.9188 0.9780 0.9475 0.9222 1.1005
CL-7B 0.8936 0.9076 0.9781 0.9415 0.8976 1.9353
QW-0.6B 0.8512 0.8774 0.9498 0.9122 0.8914 1.3213
QW-1.7B 0.8664 0.9003 0.9457 0.9225 0.8964 0.6153
QW-4B 0.8807 0.9166 0.9471 0.9316 0.9073 0.7134
QW-8B 0.8843 0.9158 0.9533 0.9342 0.9092 0.7060
Impact of Lookahead Length Selection Strategies. Both $\textsc{AdaDec}_{L=5}$ and $\textsc{AdaDec}_{dynL}$ consistently improve pass@1 accuracy over Greedy decoding across all tested models on HumanEval and MBPP benchmarks. On HumanEval, $\textsc{AdaDec}_{L=5}$ achieves a slightly higher average improvement (+4.42%) compared to $\textsc{AdaDec}_{dynL}$ (+3.96%), with $\textsc{AdaDec}_{L=5}$ outperforming $\textsc{AdaDec}_{dynL}$ on four out of eight models and achieving parity on one. Conversely, on MBPP, $\textsc{AdaDec}_{dynL}$ attains the highest average gain (+6.12%) versus $\textsc{AdaDec}_{L=5}$ (+5.69%), outperforming $\textsc{AdaDec}_{L=5}$ on four models and matching it on three. Looking at individual models, $\textsc{AdaDec}_{L=5}$ excels on DS-1.3B and QW-8B on HumanEval, showing the largest relative gains, while $\textsc{AdaDec}_{dynL}$ performs better on QW-4B and CL-7B. On MBPP, $\textsc{AdaDec}_{dynL}$ slightly edges out $\textsc{AdaDec}_{L=5}$ on larger models such as DS-6.7B and QW-8B.
Overall, $\textsc{AdaDec}_{L=5}$ tends to deliver stronger improvements on HumanEval, particularly for smaller to mid-sized models, whereas $\textsc{AdaDec}_{dynL}$ shows more consistent gains on MBPP, especially with larger models. This suggests that $\textsc{AdaDec}_{L=5}$’s fixed lookahead length may better suit certain tasks or models, while $\textsc{AdaDec}_{dynL}$’s adaptive lookahead length offers advantages on others, reflecting a tradeoff between stability and flexibility in lookahead strategy.
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Summary: The learned entropy threshold $\tau^{\bm{LM}}$ effectively distinguishes uncertain decoding steps, enabling accurate, adaptive lookahead decisions. Both $\textsc{AdaDec}_{L=5}$ and $\textsc{AdaDec}_{dynL}$ improve pass@1 over Greedy decoding, with $\textsc{AdaDec}_{L=5}$ performing better on HumanEval and $\textsc{AdaDec}_{dynL}$ showing stronger gains on MBPP, reflecting a tradeoff between fixed and dynamic lookahead strategies.
IV-E RQ4: Generalizability across Models and Datasets
We evaluate the generalizability of AdaDec along two dimensions:
Cross-Model Generalizability. We evaluate AdaDec across a broad spectrum of open-source LLMs ranging from 0.6B to 8B parameters, including code-specialized models like DeepSeek-Coder (1.3B, 6.7B), CodeLlama-7B, and StableCode-3B, as well as general-purpose models such as Qwen3 (0.6B, 1.7B, 4B, 8B). In all cases, AdaDec consistently improves Pass@1 accuracy, demonstrating strong robustness across architectures and scales.
Cross-Dataset Generalizability. We further assess AdaDec on two diverse code generation benchmarks, HumanEval and MBPP, which differ in problem complexity and code structure. As shown in Table II, AdaDec increases Pass@1 by an average of around 4% on HumanEval and around 6% on MBPP. These results confirm AdaDec’s effectiveness across a range of programming tasks.
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Summary: AdaDec generalizes well across models and datasets, consistently improving Pass@1 accuracy on LLMs from 0.6B to 8B and achieving reliable gains on both HumanEval and MBPP.
V Discussion
V-A Limitations
While AdaDec demonstrates improvements in code generation accuracy and maintains acceptable computational cost, it has several limitations that warrant further investigation:
Reliance Solely on Shannon Entropy. Our uncertainty trigger is based solely on the Shannon entropy of the token distribution, which provides a coarse measure of model uncertainty. More nuanced contextual features—such as entropy of subpopulations of tokens, token-level gradient norms, or higher-order statistics—might better capture subtle divergences. Future work could integrate richer uncertainty signals or leverage learned uncertainty estimators.
Limitations of $\textsc{AdaDec}_{dynL}$. The $\textsc{AdaDec}_{dynL}$ variant leverages explicit line breaks in Python code to determine dynamic lookahead length. While effective in this context, such structural cues may not generalize well to other domains—such as natural language generation—or to programming languages that lack clear or consistent line delimiters.
Threshold Learning Simplicity. We employ a logistic regression classifier to learn entropy thresholds, which offers interpretability but may be insufficiently expressive for capturing complex interactions among uncertainty features. More powerful models (e.g., small neural networks) could potentially improve threshold accuracy, at the cost of increased complexity.
Addressing these limitations will be an actionable direction for future work, towards making uncertainty-guided adaptive decoding more robust, flexible, and broadly applicable.
V-B Case Study
To demonstrate how AdaDec enhances decoding decisions in practice, we present a representative example from the HumanEval benchmark (Figure 5). The task is to implement the function check_dict_case, which returns True if all dictionary keys are either lowercase or uppercase, and False otherwise. The prompt provides detailed instructions along with multiple illustrative examples.
Figure 5: A case study from HumanEval
At a certain point, the model DS-1.3B has produced the initial structure shown in the “Current Seq” portion. At this stage, it must decide how to proceed. A standard greedy decoding checks islower() and isupper() directly on all keys, implicitly assuming that the keys are strings. This assumption is unsafe: if any key is a non-string (e.g., an integer), a runtime error will occur. In contrast, AdaDec detects high entropy at this decision point and triggers lookahead reranking. Among the candidate continuations, it selects a branch that first verifies whether all keys are strings before checking their case. This results in a more robust and semantically correct implementation that gracefully handles edge cases, faithfully reflecting the prompt’s intent.
This example illustrates how AdaDec uses entropy signals and semantic reranking to proactively avoid subtle bugs that are often missed by standard decoding strategies. As a result, it improves the reliability of generated code without requiring any changes to the underlying model.
VI Threats to Validity
Internal Validity. A key threat to internal validity lies in identifying drift points in our empirical study. We use a semi-automated AST structural comparison pipeline followed by human verification. While this reduces manual annotation effort, it may still miss logic-altering changes caused by identifier drift, leading to false negatives. To address this, we further conduct RQ2 in our empirical study to validate the correlation between entropy and ground-truth token ranks.
External Validity. A major threat to external validity is dataset diversity. We evaluate AdaDec on two widely used Python benchmarks—HumanEval and MBPP—which focus on function synthesis. The generalizability of our approach to other languages or larger, more complex codebases remains untested. Another potential threat is the exclusion of closed-source or commercial models. While we evaluate several open-source models (DeepSeek, StableCode, CodeLlama, Qwen3), our findings may not fully extend to proprietary models such as GPT-4 or Codex.
VII Related Work
VII-A LLM-based Code Generation
Large Language Models (LLMs) have shown strong potential in software engineering tasks such as code generation [1, 2, 3, 4, 5], unit test generation [27, 28, 29, 30, 31], and comment generation [32, 33, 34]. Typically, LLMs generate code from textual requirements via prompt-based interaction. By learning expressive representations from large-scale code and documentation corpora, they can produce syntactically correct and context-aware code snippets [8, 9, 10]. Open-foundation models further advance tasks like code completion, bug fixing [11, 12, 13], and test-case generation.
To evaluate their effectiveness, benchmarks like HumanEval [16], MBPP [26], and DS-1000 [35] focus on function-level generation, while ClassEval [36] extends evaluation to class-level tasks. More recently, CoderEval [37] and DevEval [38] target repository-level scenarios, reflecting growing research interest in applying LLMs to real-world software development.
VII-B Decoding strategies for LLMs
Large Language Models (LLMs) generate text in an autoregressive manner, producing one token at a time. At each generation step, the model samples a token from its predicted probability distribution. The most straightforward sampling strategy is greedy decoding, which selects the token with the highest probability as the next word. A more refined variant is beam search [39], which maintains multiple candidate sequences at each decoding step and preserves the top-k sequences based on cumulative log-probabilities.
Beyond these, commonly used sampling strategies include temperature sampling [40], top-p (nucleus) sampling [40], and top-k sampling [40], each designed to balance diversity and coherence in generated outputs.
To further improve sampling for code generation tasks, Li et al. introduced AdapT Sampling [41] (Adaptive Temperature Sampling). This method dynamically adjusts the temperature parameter T based on token context. For instance, when the current token initiates a code block, a higher T is used to encourage diversity and creativity. In contrast, for other tokens, a lower T is applied to suppress random noise and preserve syntactic and semantic correctness.
Another relevant approach is UnCert-CoT [42], which enhances code generation through uncertainty-aware Chain-of-Thought (CoT) reasoning, activating multi-path reasoning when token uncertainty surpasses a fixed threshold. While effective, its relies on fixed thresholds and costly CoT reasoning. In contrast, AdaDec learns model-specific uncertainty thresholds for stronger generalization, while its lightweight rerank mechanism cuts computational costs significantly without compromising benchmark performance.
VIII Conclusion
This paper presents AdaDec, an uncertainty-guided adaptive decoding framework for large language model-based code generation. Motivated by an empirical study revealing that ranking errors at high-uncertainty decoding steps often lead to logic drift, AdaDec introduces a novel pause-then-rerank mechanism that dynamically identifies and corrects uncertain token predictions using Shannon entropy. By learning model-specific entropy thresholds and employing a lookahead-based reranking strategy, AdaDec achieves substantial improvements in both accuracy and efficiency. Experiments across HumanEval and MBPP benchmarks demonstrate that AdaDec consistently outperforms greedy decoding and competes favorably with beam search while significantly reducing decoding overhead. These findings underscore the importance of token-level uncertainty modeling and open up new directions for adaptive decoding in code generation and beyond.