### v3

# Towards Better Code Generation: Adaptive Decoding with Uncertainty Guidance

Abstract.
Code generation with large language models (LLMs) is highly sensitive to token selection during decoding, particularly at decision points where uncertainty strongly affects program correctness. Conventional strategies such as greedy decoding treat all tokens uniformly and fail to capture the uncertainty characteristics unique to code, often resulting in suboptimal outputs. In this work, we conduct an empirical analysis and show that a large fraction of generation errors arises from token misranking at high-uncertainty positions, where the correct token is available but not prioritized.
To address this, we introduce AdaDec, an adaptive decoding framework that employs a lookahead-based, uncertainty-aware pause-and-rerank mechanism. AdaDec automatically learns model-specific uncertainty thresholds and selectively invokes reranking when high uncertainty is detected, leveraging lookahead to refine token choice. Across HumanEval+, MBPP+, and DevEval benchmarks, AdaDec yields substantial improvements, achieving up to 20.9% absolute gains in Pass@1 accuracy compared with greedy decoding, while consistently outperforming prior adaptive decoding approaches such as AdapT. Furthermore, by applying reranking only when necessary, AdaDec reduces computational overhead and latency, enhancing efficiency alongside reliability. These findings underscore the value of uncertainty-guided decoding strategies in advancing the robustness and practicality of LLM-based code generation.
Code Generation, Adaptive Strategy, Language Model
1. Introduction
Large language models (LLMs) have significantly advanced the field of code generation by learning expressive representations of programming syntax and semantics from large-scale corpora of source code and documentation (Li et al., 2023; Di and Zhang, 2025; Tian et al., 2024; Jiang et al., 2025; Lin et al., 2024). Models such as CodeLlama (Rozière et al., 2024) and DeepSeek-Coder (Guo et al., 2024)—pre-trained on billions of lines of code—exhibit strong capabilities in producing syntactically correct and context-aware code snippets from natural language prompts (Yuan et al., 2023; Sun et al., 2024; Ugare et al., 2024). These capabilities have catalyzed a range of intelligent software engineering tasks, including code completion, automated bug fixing (Fu et al., 2022; Jin et al., 2023; Xia and Zhang, 2022), and test-case generation, ultimately reducing developer effort and accelerating the software development process (Jiang et al., 2024; Huynh and Lin, 2025).
At the core of code generation with large language models (LLMs) lies the decoding process, which translates learned latent representations into discrete, executable token sequences (Chen et al., 2021a). Decoding for code generation, however, faces domain-specific challenges. Code exhibits distinctive patterns of token uncertainty: certain tokens (e.g., identifiers or control keywords at the start of a line) are intrinsically ambiguous because their correct choice depends on future logic, while others (e.g., syntactically constrained keywords or repeated identifiers) are relatively deterministic. Crucially, mistakes at a few high-uncertainty points can yield irreversible semantic errors (for example, selecting “for” versus “return” at line start fundamentally alters control flow). Figure 1 provides an example. Therefore, the choice of decoding strategy—how the next token is selected or processed—plays a decisive role in balancing efficiency, diversity, and reliability.
Figure 1. Illustration of irreversible semantic errors in code due to high-uncertainty points.
A range of strategies have been proposed: deterministic greedy decoding (Radford et al., 2019) is simple and efficient but prone to premature commitment to suboptimal tokens; sampling-based approaches such as top-$k$ (Holtzman et al., 2020) and nucleus (top-$p$) (Holtzman et al., 2020) sampling enhance diversity but often introduce syntactic or semantic errors that are unacceptable for code; and uncertainty-based methods aim to exploit token-level difficulty signals. For instance, AdapT (Zhu et al., 2023) estimates uncertainty by computing cross-entropy loss between model predictions and ground truth tokens, finding that tokens with higher loss values tend to cluster at line starts. Based on this pre-identified pattern rather than dynamic signals during decoding, AdapT assigns higher temperatures to line-initial tokens, but its reliance on randomness makes gains unstable and limited. By contrast, Uncert-CoT (Zhu et al., 2025) dynamically measures uncertainty at decoding time, employ Shannon entropy (Shannon, 1948) or probability difference (the gap between the top-1 and top-2 token probabilities) to decide whether a line-initial token is highly uncertain, in which case single-line chain-of-thought reasoning is triggered. While this avoids uniform treatment of all line starts, it overlooks uncertain tokens beyond line boundaries and has shown only marginal improvements in practice. Overall, existing methods either treat all decoding steps uniformly or rely on coarse heuristics tied to line structure, and current uncertainty-based decoding approaches have not yet explored which uncertainty signals are most appropriate to guide intervention at decoding time.
To address this gap, we conduct an empirical study to better understand token-level uncertainty during the decoding process in LLM-based code generation. Using DeepSeek-Coder (Guo et al., 2024) and the HumanEval dataset (Chen et al., 2021b), we analyze decoding failures by comparing model predictions that fail test cases with their corresponding ground-truth solutions. Finding-1: The analysis reveals that many errors stem from local ranking mistakes—where the correct token exists in the candidate set but is not ranked first—thereby causing logic drift. To further investigate this phenomenon, we perform next-token prediction on 164 HumanEval problems, recording decoding information and the rank of the ground-truth token at each step. This enables us to statistically identify the locations of logic drift points and analyze the relationship between logic drift and entropy. Finding-2: The results suggest that entropy can serve as a useful signal for detecting logic drift and guiding adaptive pausing and reranking. However, due to variability across models, a fixed entropy threshold is suboptimal, highlighting the need for a dynamic, model-aware thresholding mechanism.
Base on these findings, We introduce AdaDec, an uncertainty-guided adaptive decoding framework for code generation with large language models. AdaDec enhances standard decoding (e.g., greedy decoding, sampling) by incorporating a pause-then-rerank mechanism that reacts to model uncertainty, measured by Shannon entropy at each step. When uncertainty exceeds a threshold, decoding is paused and a reranking procedure is triggered. Because a fixed threshold does not generalize well across different models, AdaDec learns this threshold through a data-driven logistic regression approach. During reranking, AdaDec selects the top candidate tokens and evaluates them using a lookahead strategy inspired by A* search heuristics and a prior constrained generation method (Lu et al., 2021), scoring each candidate based on expected future continuations. The highest-scoring token is chosen for generation, and the process repeats until the sequence is complete. In contrast to prior approaches that either treat all decoding steps uniformly or rely on coarse heuristics tied to line structure, AdaDec systematically explores Shannon entropy as a token-level uncertainty signal and applies dynamic, token-level intervention across both line-initial and in-line tokens only at high-uncertainty steps, addressing the identified gap and enables targeted correction of the most consequential decision points.
We conduct extensive experiments on the HumanEval+, MBPP+ and DevEval benchmarks using models ranging from 0.6B to 8B parameters. Our uncertainty-guided adaptive decoding framework, AdaDec, consistently improves Pass1 accuracy by up to 20.9% over Greedy decoding and consistently outperforms state-of-the-art adaptive decoding methods such as AdapT. In practice, AdaDec selectively pauses decoding at a low rate (averaging 7.59%), and the targeted lookahead introduces moderate additional latency, effectively balancing quality and efficiency. Ablation studies confirm that the learned entropy threshold accurately identifies uncertain decoding steps, and empirical comparisons show that using the learned threshold outperforms a fixed threshold, particularly for models where the fixed threshold deviates substantially from the learned value. Overall, AdaDec generalizes well across model scales and datasets, significantly improving code generation accuracy without incurring excessive runtime overhead.
In summary, this paper makes the following contributions:
  • An empirical study that reveals many code generation errors arise from local token ranking mistakes, not only occurring at line beginnings but also sometimes within lines, and that Shannon entropy serves as a useful signal for detecting uncertain decoding steps prone to logic drift.
  • An adaptive decoding framework, AdaDec, which employs an entropy-guided pause-then-rerank mechanism with learned, model-specific thresholds and a lookahead strategy to improve generation quality.
  • Experimental results that show AdaDec improves Pass@1 accuracy by up to 20.9% over Greedy decoding, achieves high efficiency by selectively pausing decoding at a low rate, and generalizes effectively across diverse model sizes and benchmark datasets.
2. Empirical Study
We conduct an empirical study to investigate the characteristics of token uncertainty in LLM-based code generation and analyze the feasibility of integrating token uncertainty into decoding process. Specifically, the study aims to answer the following research questions:
  • RQ1: Where do logic drifts occur in code, and how are they influenced by token uncertainty?
  • RQ2: How can token uncertainty be leveraged to improve token selection during decoding?
2.1. Study Setup
2.1.1. Models
To ensure the generality of our findings, we evaluate AdaDec on a diverse set of open-source code generation models with varying scales and architectures. Specifically, we include:
  • DeepSeek-Coder-1.3B-Instruct (DS-1.3B), DeepSeek-Coder-6.7B-Instruct (DS-6.7B) (Guo et al., 2024): These models are trained from scratch on 2 trillion tokens comprising 87% code and 13% natural language, employing a fill-in-the-blank objective and a 16K context window to support project-level code completion and infilling tasks.
  • StableCode-Instruct-3B (ST-3B) (Pinnaparaju et al., 2023): A 2.7B parameter decoder-only model fine-tuned using Direct Preference Optimization (DPO), demonstrating state-of-the-art performance among similarly sized models on MultiPL-E and MT-Bench coding benchmarks.
  • Qwen3-0.6B (QW-0.6B), Qwen3-1.7B (QW-1.7B), Qwen3-4B (QW-4B), Qwen3-8B (QW-8B) (Team, 2025): A family of multilingual large language models ranging from 0.6B to 8B parameters, featuring dynamic mode switching between reasoning and fast-response modes, and supporting a wide range of tasks including code generation.
These models span from 0.6B to 8B parameters and are accessed via HuggingFace Transformers. They cover a range of architectures and instruction-tuning schemes, ensuring a diverse evaluation across different model families and sizes.
2.1.2. Implementation
Our implementation is based on PyTorch and HuggingFace Transformers, and decoding is performed in a batch-free, single-example fashion to avoid interference between instances. All models are run using half-precision (i.e., torch.float16) to accelerate inference and reduce memory usage. The maximum number of generated tokens is capped at 512 for all methods to ensure consistent comparison across models.
Experiments are conducted on a machine equipped with NVIDIA RTX 5880 Ada Generation GPUs (48 GB VRAM each) and an NVIDIA A100 GPU (80 GB VRAM), supported by NVIDIA Driver 550.163.01 and CUDA 12.4.
2.2. RQ1: Where do logic drifts occur in code, and how are they influenced by token uncertainty?
To address this research question, we first use DS-6.7B as a representative model to generate code for 164 programming problems in the widely used HumanEval dataset (Chen et al., 2021b). We begin by analyzing the locations of logic drifts—points during decoding where the generated sequence diverges from the correct solution—which often lead to incorrect program behavior. We then investigate how token-level uncertainty, quantified using Shannon entropy (Shannon, 1948), relates to the occurrence of such drifts.
2.2.1. Method
For each programming problem in the HumanEval dataset, we use DS-6.7B to generate code snippets using a greedy decoding strategy. We then filter out the generations that successfully pass all test cases. Based on an initial human analysis, we observe that in several cases, the model begins with a correct sequence but introduces a critical error—often in the form of incorrect code structures—at a specific step. At these error points, the correct next token is not missing from the candidate set but is ranked below the top-1 position (commonly second or third), resulting in an irreversible deviation from the correct logic. This suggests that some generation failures are caused by local ranking errors at specific decoding steps, ultimately leading to incorrect code behavior. We refer to such steps as logic drift points during decoding and aim to identify them.
Given the high cost of fully manual annotation, we develop a semi-automatic approach that combines abstract syntax tree (AST) comparison with human verification and fine-grained token-level annotation. To ensure annotation focuses only on faulty generations, we first filter out all generated samples that can pass the tests. Subsequently, we construct ASTs for both the ground-truth and generated code, and normalize them by removing identifiers and comments. This normalization ensures that the comparison focuses solely on structural differences. We then perform a parallel traversal to compare the two normalized ASTs node by node. A mismatch is flagged when node types differ, operators or constants are inconsistent, or discrepancies occur in control structures. The mismatch is then traced back to its corresponding line in the original code and extracted as a candidate drift point. For each mismatched line, we further determine whether the drift occurs at the beginning of the line and record this information for statistical analysis. Finally, each candidate drift point is reviewed by a human annotator to confirm whether it represents an actual drift point and to annotate the precise token where the drift begins. For example, a mismatch between the generated line if shift > len(s): and the ground-truth line if shift >= len(digits): triggers a drift point at the position of the operator >=.
In addition to the identified drift points, we collect all other non-drift decoding steps from the same generated code snippets. We compute the Shannon entropy for each step and compare the entropy distribution of drift points with that of non-drift steps. This comparison allows us to investigate the relationship between logic drift and token uncertainty, as measured by Shannon entropy. Given a probability distribution $\bm{p}$, the corresponding entropy $H(\bm{p})$ is calculated as follows:
$H(\bm{p})=-\sum_{p_{i}\in\bm{p}}p_{i}\log p_{i}$
2.2.2. Results
We identify a total of 46 drift points and 2,889 non-drift decoding steps across 46 HumanEval problems. Statistical analysis shows that divergences at line start account for 86.96% of all drift points, while 13.04% occur within lines. This observation suggests that focusing solely on line-initial tokens for uncertainty-based intervention is insufficient, as it ignores more than 10% of logic drift points. Figure 2 presents a boxplot comparing the entropy distribution at drift points with that of all non-drift steps. For clarity, outliers are omitted in boxplots. The entropy values at drift points are markedly higher: the median entropy is 1.26, with a mean of 1.39, compared to a much lower median of 0.03 and mean of 0.30 for the rest. The interquartile range (IQR) for drift points is also shifted higher (0.63–2.01) relative to that of non-drift steps (0.003–0.34). The results show that tokens at drift points exhibit significantly higher entropy values compared to the rest, suggesting a clear relationship between entropy and logic drift.
Figure 2. Entropy comparison between drift points and non-drift decoding steps.
Finding 1: This empirical pattern indicates that focusing solely on line-start tokens for uncertainty-based intervention is insufficient, motivating the search for a suitable uncertainty signal that monitors all tokens and triggers targeted intervention at high-uncertainty steps. Subsequent analysis shows that token uncertainty measured by Shannon entropy is strongly correlated with logic drift, providing a useful signal for identifying potential drift points and predicting code behavior deviations during decoding.
2.3. RQ2: How can token uncertainty be leveraged to improve token selection during decoding?
To integrate token uncertainty into the token selection process and enhance the likelihood of generating correct code, we further investigate the relationship between entropy and the rank positions of expected tokens during decoding.
2.3.1. Method
We perform next-token prediction using 164 problems and their corresponding ground-truth solutions from the HumanEval dataset. Specifically, given a problem $P$ and its ground-truth solution $S$, we tokenize them into sequences $\mathrm{T}^{P}$ and $\mathrm{T}^{S}$, respectively. At each decoding step $k$, we feed the concatenated sequence $\mathrm{T}^{P}\oplus\mathrm{T}^{S}_{:k}$ into the LLM to obtain a probability distribution over the next token. Here, $\mathrm{T}^{S}_{:k}$ denotes the prefix of the ground-truth solution containing the first $k$ tokens, and $\oplus$ indicates sequence concatenation. By iterating over all positions $k$ from 1 to the length of $\mathrm{T}^{S}$, we collect a set of predicted probability distributions. For each prediction, we compute the Shannon entropy of the distribution and record the rank of the expected (ground-truth) token within it.
We apply this next-token prediction procedure to a range of open-source LLMs, including three code-specific models, DS-1.3B, DS-6.7B, and ST-3B, as well as four general-purpose models from the Qwen3 family: QW-0.6B, QW-1.7B, QW-4B, and QW-8B. For Qwen3, we disable the reasoning-specific settings (e.g., thinking mode). In total, we collect 180,879 token-level probability distributions for two DeepSeek-Coder models, 169,496 for StableCode, and 129,165 for Qwen3 series.
2.3.2. Results
We compute the Spearman correlation between entropy and the rank of the ground-truth token across all models. As shown in Table 1, the results consistently exhibit a positive correlation, further reinforcing the notion that entropy serves as a meaningful signal for identifying uncertain decoding steps. We omit p-values from the table, as the large sample sizes (approximately 130,000 to 180,000 tokens per model) yield statistically significant results with p-values effectively indistinguishable from zero.
Table 1. Spearman correlation $\bm{\rho}$ between entropy and the rank of ground-truth tokens.
Model DS-1.3B DS-6.7B ST-3B QW-0.6B QW-1.7B QW-4B QW-8B
$\bm{\rho}$ 0.4728 0.4595 0.4620 0.5249 0.5028 0.4936 0.4877
Figure 4 shows how the percentage of decoding steps with entropy exceeding a given threshold changes as the threshold increases. Figure 4 illustrates the corresponding change in the average rank of ground-truth tokens, separated by whether the entropy at each step is above or below the threshold. These plots reveal that different models exhibit varying sensitivity to entropy thresholds. For example, at a threshold of 1, QW-0.6B has approximately 15% of decoding steps exceeding the threshold, while other models remain below 10%. In terms of average rank, two models (DS-6.7B and QW-1.7B) exceed a value of 4.5, and the rest fall in the range of 3.5 to 4.
Figure 3. Change in the average rank of ground-truth tokens above and below a given entropy threshold, as the threshold increases.
Figure 4. Change in the percentage of decoding steps exceeding a given entropy threshold as the threshold increases.
Finding 2: The observed correlation between entropy and the rank of the ground-truth token suggests that entropy can be used as an indicator to adaptively pause the decoding process and rerank uncertain tokens. However, our entropy percentile analysis shows that it is difficult to define a universal, fixed entropy threshold across all models that effectively balances pause frequency and the number of reranking candidates. This highlights the need for a dynamic mechanism to determine the model-specific entropy threshold.
3. Approach
We now introduce AdaDec, an uncertainty-guided adaptive decoding framework for LLM-based code generation. AdaDec integrates a token-level pause-then-rerank mechanism into the decoding process of the target LLM, drawing on insights from our empirical analysis.
3.1. Overview
Figure 5. Approach Overview of AdaDec
As illustrated in Figure 5, AdaDec is an adaptive (dynamic) decoding framework: at inference time it continuously monitors per-step uncertainty and dynamically decides whether to intervene, rather than applying a uniform policy across all steps or only at line-start positions. Given a target model LM, at each decoding step $t$ the model produces a next-token distribution $\bm{p}_{t}$. AdaDec computes the Shannon entropy $H(\bm{p}_{t})$ and compares it against a model-specific threshold $\tau^{\bm{LM}}$. If $H(\bm{p}_{t})\leq\tau^{\bm{LM}}$, the framework proceeds with the base decoding policy (e.g., greedy, top-$k$, or top-$p$ sampling). If $H(\bm{p}_{t})>\tau^{\bm{LM}}$, AdaDec pauses normal decoding and triggers a targeted reranking routine over the top-$K$ candidates. During reranking, each candidate undergoes a short lookahead of up to $N$ steps to estimate the quality of its future trajectory, and the candidate with the highest score is selected as the next token. This per-step decision loop repeats until an end-of-sequence (EOS) token is generated.
The adaptive nature of AdaDec has three practical consequences. First, the intervention is selective and data-driven: only high-uncertainty steps incur additional computation, which limits overhead compared to always-on strategies. Second, the threshold $\tau^{\bm{LM}}$ is learned in a preparation stage (we describe the learning procedure in Sec. 3.2) so that the pause decision is model-aware rather than hand-tuned. Third, the framework supports flexible reranking configurations (e.g., adjustable lookahead length $N$ and lookahead width $K$), allowing trade-offs between accuracy and latency to be tuned for different deployment constraints.
AdaDec is designed to be broadly applicable. Since the pause-then-rerank mechanism is triggered purely by uncertainty signals (entropy), it can be seamlessly integrated with various decoding strategies, including sampling. This flexibility underscores the generality of AdaDec: rather than being tied to a specific decoding style, it acts as a plug-in framework that can enhance both deterministic (e.g., greedy) and stochastic (e.g., sampling) generation. Unless otherwise specified, we use greedy decoding as the base policy throughout the remainder of this paper.
The details of the pause checking and lookahead reranking are described in the following sections.
3.2. Entropy‐based Pause Checking
We use the Shannon entropy $H(\bm{p}_{t})$ over the probability distribution $\bm{p}_{t}$ predicted by the model LM to measure the token uncertainty. When $H$ exceeds a threshold, AdaDec pauses the decoding process. As revealed in the empirical study, a fixed entropy threshold may not generalize well across models with different output distributions. Therefore, we propose a data-driven method to learn a model-specific threshold $\tau^{\bm{LM}}$ using logistic regression.
3.2.1. Data Collection
To enable threshold learning, we cast the problem as a binary classification task. Each sample corresponds to a decoding step and is labeled based on whether the ground truth token is ranked top-1 in the predicted distribution. The input feature for classification is the entropy $H(\bm{p}_{t})$, which reflects the model’s uncertainty at that step.
We generate training data by conducting ground-truth–guided next-token prediction with the target model LM on the BigCodeBench dataset (Zhuo et al., 2024), while carefully ensuring that the construction of our training data does not overlap with the evaluation benchmarks (HumanEval+, MBPP+, and DevEval+). To mitigate the risk of data contamination, we manually reviewed problem statements and function signatures to check for potential duplicates or near-duplicates between BigCodeBench and the evaluation sets. In this process, the ground truth token is provided as input at each step instead of the model’s own prediction. During this process, we record the model’s predicted probability distribution over the vocabulary at each decoding step and calculate the entropy of the predicted distribution. Based on whether the ground truth token is ranked as the top-1 prediction, we assign a binary label: 1 if it is top-1, and 0 otherwise. Based on this process, we collect between approximately 100,000 and 160,000 positive samples (labeled as 1) and between 17,000 and 23,000 negative samples (labeled as 0) per model. For example, in the case of DS-6.7B, we obtain 161,856 positive samples and 19,023 negative samples.
We then balance the dataset by randomly downsampling the positive samples and selecting an equal number of negative samples, which is a common technique for handling imbalanced classification tasks (Lemaitre et al., 2016). This choice is driven by the nature of logic drift in code generation: once a drift point occurs, it typically leads to irreversible semantic errors. Therefore, it is essential to maximize the classifier’s ability to detect negative cases, even if they are relatively scarce. By constructing a balanced dataset, we prevent the model from being biased toward the majority class and enable it to more effectively learn the decision boundary for identifying uncertain decoding steps. This strategy enhances the classifier’s sensitivity to potential drift points, thereby improving the overall reliability of the entropy-based triggering mechanism.
3.2.2. Regression Model Training
We train a logistic regression model to predict whether the ground truth token is the top-1 prediction, using only the entropy value as input. The model estimates the probability of a positive label using the sigmoid function:
$P(y=1\mid H^{(i)})=\frac{1}{1+\exp(-(\beta_{0}+\beta_{1}H^{(i)}))},$
where $H^{(i)}$ denotes the entropy of the $i$-th sample, and $\beta_{0}$, $\beta_{1}$ are the model parameters to be learned. These parameters are learned by minimizing the binary cross-entropy loss through maximum likelihood estimation. Note that, since the training data is collected using the target LM, the trained logistic regression model is specific to that LM.
3.2.3. Threshold Estimation
To estimate a model-specific entropy threshold $\tau^{\bm{LM}}$, we use the trained logistic regression model to perform binary classification on the validation set, where the task is to predict whether the top-1 token matches the ground-truth token. As the logistic regression model outputs probabilities, we first identify the optimal probability threshold that maximizes classification accuracy, and then convert this probability threshold into an equivalent entropy threshold.
Specifically, we conduct a grid search over probability thresholds in the range [0.01, 0.99], using a step size of 0.01. For each candidate threshold, a sample is classified as positive if its predicted probability exceeds the threshold, and negative otherwise. The threshold that achieves the highest classification accuracy on the validation set is selected as the optimal probability threshold, denoted by $p^{*}$. We then convert this optimal probability threshold into the corresponding entropy threshold using the parameters of the trained logistic regression model. Given the intercept $\beta_{0}$ and the coefficient $\beta_{1}$, the entropy threshold $\tau^{\bm{LM}}$ is computed using the log-odds transformation:
$\tau^{\bm{LM}}=\frac{\log\left(\frac{p^{*}}{1-p^{*}}\right)-\beta_{0}}{\beta_{1}}.$
Entropy values below $\tau^{\bm{LM}}$ are classified as positive (i.e., the model is likely to correctly predict the top-1 token), while those above are classified as negative. This entropy threshold is then used as the final decision boundary for the target language model during decoding.
3.2.4. Pause Triggering
During decoding, AdaDec uses the learned threshold $\tau^{\bm{LM}}$ to decide when to trigger candidate reranking. If the entropy at a decoding step exceeds $\tau^{\bm{LM}}$, it indicates high uncertainty, suggesting that the top-1 prediction may be incorrect. In such cases, the decoding process is paused, and the top candidates are reranked to improve accuracy.
3.3. Lookahead-based Token Reranking
AdaDec employs a lookahead re-scoring strategy for token reranking, enhancing selection by assessing the future impact of each candidate token at every decoding step. Inspired by A* search heuristics and prior work on constrained generation (Lu et al., 2021), this method extends token probability evaluation beyond immediate next-token predictions to consider longer-term sequence quality. The process is governed by two hyperparameters: the lookahead beam size ($B$), which determines the number of candidate tokens considered, and the lookahead length ($L$), which sets the length of the future token sequence evaluated.
3.3.1. Lookahead Process
At decoding step $t$, the model outputs a probability distribution $\bm{p}_{t}$ over the vocabulary for the next token. From this distribution, we select the top-$B$ tokens to form the candidate set $\mathcal{C}_{t}=\{c_{t}^{(1)},c_{t}^{(2)},\dots,c_{t}^{(B)}\}$.
For each candidate token $c_{t}^{(k)}$, we construct a lookahead trajectory of length $L$ by appending $L$ tokens generated via greedy decoding, conditioned on the current input and previously generated tokens. This results in the trajectory $\mathbf{T}_{t}^{(k)}$:
$\mathbf{T}_{t}^{(k)}=(c_{t}^{(k)},\hat{c}_{t+1}^{(k)},\hat{c}_{t+2}^{(k)},\dots,\hat{c}_{t+L}^{(k)}),$
where each $\hat{c}_{t+i}^{(k)}$ is the most probable token predicted at step $t+i$, given the prefix ending with $c_{t}^{(k)}$.
Trajectory Scoring. We evaluate each trajectory by computing the average log-probability of its tokens:
$\text{score}(c_{t}^{(k)})=\frac{1}{L+1}\left[\log P(c_{t}^{(k)})+\sum_{i=1}^{L}\log P(\hat{c}_{t+i}^{(k)})\right],$
where $P(\cdot)$ denotes the model’s predicted probability for a token. Because we apply an early stopping mechanism, different candidates may result in varying lookahead lengths. To ensure a fair comparison across such candidates, we normalize the total log-probability by computing the geometric mean of token probabilities in the above equation, rather than relying solely on log-likelihood. This normalization mitigates the bias against longer sequences, which tend to accumulate lower joint probabilities. The resulting score captures the model’s overall confidence in the full trajectory, balancing immediate token likelihood with the coherence of future predictions.
Token Selection. The candidate with the highest trajectory score is selected as the output token at step $t$:
$c^{*}_{t}=\arg\max_{c_{t}^{(k)}\in\mathcal{C}_{t}}\text{score}(c_{t}^{(k)}).$
This process is repeated at each decoding step, ensuring that token selection is guided not just by immediate probabilities, but by the expected quality of the downstream sequence.
This lookahead mechanism improves generation quality by favoring tokens that lead to fluent and coherent continuations, which is particularly important in structured domains like code generation. The hyperparameters $B$ and $L$ control the trade-off between computational cost and generation quality: larger values offer better foresight at the expense of increased inference time.
3.3.2. Lookahead Beam Size Selection
To determine an appropriate value for $B$, we analyze the training data collected in Section 3.2.1. Specifically, we examine the average rank of the ground truth token at decoding steps where the Shannon entropy is either above or below the threshold $\tau^{\bm{LM}}$, which is learned via logistic regression. After filtering out samples where the ground truth token’s rank exceeds 20, we observe a clear distinction: when the entropy is below or equal to the threshold, the average rank typically falls between 1.1 and 1.3, indicating high model confidence. In contrast, when the entropy exceeds the threshold—thus triggering a pause—the average rank increases to between 2.6 and 3.4.
These findings support the use of a small beam size (e.g., $B=3$) to achieve a practical trade-off between candidate coverage and computational efficiency.
3.3.3. Lookahead Length Selection
We adopt a lookahead length parameter $L$ when performing trajectory scoring. At each reranking step we simulate up to $L$ future tokens for each candidate to estimate the quality of its continuation. The parameter $L$ controls the trade-off between accuracy and efficiency: a larger $L$ provides a more reliable estimate of long-term trajectory quality but incurs higher computation cost, whereas a smaller $L$ reduces overhead but may fail to capture meaningful differences among candidates.
In practice, we find that moderate values of $L$ (e.g., $L=5$) are sufficient to balance these concerns, since code generation typically exhibits strong local dependencies and short-horizon lookahead already provides useful signals. We analyze the effect of varying $L$ through experiments in Section 4.5.
4. Evaluation
We conduct extensive experiments to evaluate the effectiveness of AdaDec. Specifically, we aim to answer the following research questions:
  • RQ1 (Effectiveness): How effective is AdaDec in improving the functional correctness of generated code?
  • RQ2 (Efficiency): How efficient is AdaDec in terms of computational cost?
  • RQ3 (Ablation Study): How effective is the logistic regression–based threshold learning component, and can the learned thresholds better guide practical decoding?
  • RQ4 (Parameter Sensitivity): How does AdaDec perform under different values of the lookahead length $L$?
4.1. Experimental Setup
Our experimental setup is as follows.
Datasets. We evaluate our method on three widely used code generation benchmarks:
  • HumanEval+ (Liu et al., 2023): An extended version of the original HumanEval(Chen et al., 2021b) dataset, comprising 164 hand-crafted Python programming problems. The number of test cases has been increased by nearly 81×, enabling a more comprehensive evaluation of the functional correctness of generated code.
  • MBPP+ (Liu et al., 2023): A refined and optimized subset of the original MBPP(Austin et al., 2021) dataset, comprising 378 Python programming problems. MBPP+ reduces the number of problems compared to MBPP but increases the size of the unit tests by 35×, providing a much stricter evaluation of the functional correctness of LLM-generated code.
  • DevEval (Li et al., 2024): A repository-level benchmark aligned with real-world code repositories. It includes 1,874 testing samples from 117 repositories across 10 domains, providing a comprehensive evaluation of code generation models in practical software development scenarios. In our experiments, we randomly sample 238 problems. This sample size follows the statistical guideline that 238 measurements are sufficient to ensure, with 90% confidence, that the estimated performance lies within a ±5% margin of the true value.
Baselines. We compare AdaDec against two standard decoding strategies:
  • Greedy Decoding (Radford et al., 2019): A straightforward approach where at each step, the model selects the token with the highest probability as the next token. This method is computationally efficient but may lead to suboptimal solutions due to its myopic nature.
  • AdapT (Zhu et al., 2023): An uncertainty-guided adaptive decoding strategy that adjusts the sampling temperature for tokens at the beginning of a line (high-uncertainty tokens) to increase randomness and help overcome difficult-to-predict points, while keeping low-uncertainty tokens more deterministic. For our experiments, we set the parameters as follows: $a=0.5$ for high-uncertainty tokens and $b=0.05$ for all other tokens.
We also considered Uncert-CoT (Zhu et al., 2025), another uncertainty-guided decoding method originally proposed for chain-of-thought reasoning tasks. However, we do not include it in our experimental comparison since no official implementation is available and the original paper omits several crucial details necessary for faithful reproduction.
4.2. RQ1: Effectiveness in Improving Functional Correctness
Table 2 presents the Pass@1 performance of three decoding strategies, Greedy, AdapT, and AdaDec, across seven LLMs evaluated on the HumanEval+, MBPP+ and DevEval datasets.
Table 2. Pass@1 comparison of Greedy, AdapT, and AdaDec on HumanEval+, MBPP+, and DevEval. Values in parentheses indicate the change relative to Greedy.
Model HumanEval+ MBPP+ DevEval
Greedy AdapT AdaDec Greedy AdapT AdaDec Greedy AdapT AdaDec
DS-1.3B 51.83% 52.44%(+0.61%) 57.32%(+5.49%) 41.01% 41.27%(+0.26%) 49.47%(+8.46%) 13.45% 13.03%(-0.42%) 18.49%(+5.04%)
DS-6.7B 62.80% 63.41%(+0.61%) 65.24%(+2.44%) 58.47% 58.73%(+0.26%) 61.11%(+2.64%) 28.57% 28.57%(0.00%) 27.73%(-0.84%)
ST-3B 48.78% 51.83%(+3.05%) 50.61%(+1.83%) 38.89% 39.42%(+0.53%) 46.03%(+7.14%) 15.55% 13.03%(-2.52%) 22.69%(+7.14%)
QW-0.6B 17.07% 18.29%(+1.22%) 22.56%(+5.49%) 9.52% 11.11%(+1.59%) 30.42%(+20.90%) 2.94% 2.52%(-0.42%) 7.56%(+4.62%)
QW-1.7B 40.24% 42.07%(+1.83%) 44.51%(+4.27%) 37.57% 36.51%(-1.06%) 42.86%(+5.29%) 7.98% 7.56%(-0.42%) 11.76%(+3.78%)
QW-4B 53.05% 53.66%(+0.61%) 56.10%(+3.05%) 40.48% 41.01%(+0.53%) 51.32%(+10.84%) 9.24% 9.24%(0.0%) 16.81%(+7.57%)
QW-8B 56.71% 57.32%(+0.61%) 62.80%(+6.09%) 29.37% 29.63%(+0.26%) 47.09%(+17.72%) 4.62% 4.20%(-0.42%) 7.56%(+2.94%)
AVG. $\Delta$ +1.22% +4.09% +0.34% +10.43% -0.60% +4.32%
Comparison with Greedy Search. Across all three benchmarks, AdaDec consistently outperforms Greedy decoding across most datasets and models. On HumanEval+, AdaDec improves pass@1 by an average of +4.09%, with gains as high as +6.09% for QW-8B and +5.49% for DS-1.3B. On MBPP+, the improvement is even more pronounced: AdaDec yields an average gain of +10.43%, with the most substantial boost observed for QW-0.6B (+20.90%) and QW-8B (+17.72%). On DevEval, AdaDec also yields meaningful improvements for most models (+4.32% on average), with ST-3B and QW-4B showing particularly strong gains (+7.14% and +7.57%, respectively). The only exception is DS-6.7B, where we observe a slight drop ($0.84\%$). This may be due to DevEval’s repository-level nature, where future logic are harder to capture with short-horizon lookahead, and DS-6.7B already performs strongly under Greedy decoding, leaving less room for improvement.
Comparison with AdapT. When compared to AdapT, AdaDec again shows clear advantages. While AdapT offers marginal improvements over Greedy decoding (e.g., +1.22% on HumanEval+ and +0.34% on MBPP+), it sometimes underperforms (e.g., -0.60% on DevEval). In contrast, AdaDec consistently achieves larger gains, surpassing AdapT by wide margins across nearly all models and datasets. For example, on MBPP+, AdaDec exceeds AdapT by more than +10% on average, with the gap reaching +19.31% for QW-0.6B. On HumanEval+, AdaDec achieves substantially higher gains (average +4.09% vs. +1.22%), highlighting its stronger effectiveness in enhancing code correctness.
Summary: Overall, the results indicate that AdaDec is highly effective in improving the functional correctness of generated code. Compared with Greedy search, it achieves consistent and often substantial improvements across most models and benchmarks. Compared with AdapT, AdaDec demonstrates more robust and scalable gains. These findings confirm the effectiveness of the proposed approach in enhancing decoding quality.
4.3. RQ2: Efficiency in Computational Cost
We evaluate the efficiency of AdaDec in terms of computational cost on the HumanEval+, MBPP+ and DevEval datasets from the following perspectives. First, we measure the pause rate of AdaDec to assess the possible overhead introduced by its pause-then-rank mechanism. Second, we compare the prediction latency of AdaDec against Greedy decoding and AdapT method.
Pause Rate. As presented in Table 3, the pause rate ranges from 4.80% to 12.17%, with an average of 7.59%. This low activation frequency indicates that AdaDec concentrates computational effort on a small subset of high-uncertainty steps, effectively balancing local correction with overall efficiency. Across all datasets, the pause rate remains below 10% for all models except QW-1.7B, demonstrating that selective triggering successfully avoids the high cost associated with global lookahead.
Table 3. Pause ratio of the models on HumanEval+, MBPP+, and DevEval.
DS-1.3B DS-6.7B ST-3B QW-0.6B QW-1.7B QW-4B QW-8B
HumanEval+ 6.51% 5.48% 4.80% 6.11% 11.17% 7.82% 5.98%
MBPP+ 8.90% 7.46% 6.52% 8.50% 12.17% 7.11% 8.33%
DevEval 7.41% 6.17% 6.59% 6.16% 11.47% 9.69% 5.04%
AVG. $\Delta$ 7.61% 6.37% 5.97% 6.92% 11.60% 8.21% 6.45%
Latency. Table 4 reports the average generation time per problem for Greedy, AdapT, and AdaDec across HumanEval+, MBPP+, and DevEval. Overall, AdaDec introduces moderate additional latency compared to Greedy, reflecting the computational overhead of the pause-and-rerank mechanism. On most models and datasets, the increase remains reasonable—for instance, +0.23s on HumanEval+ and -0.84s on MBPP+ on average—demonstrating that selective triggering effectively balances enhanced generation quality with efficiency.
AdapT exhibits mixed runtime behavior. In some cases, it is faster than Greedy (e.g., DS-1.3B on MBPP+), likely because higher-temperature sampling applied only at high-uncertainty tokens (typically line-start tokens) can lead to shorter, simpler continuations, reducing the total number of decoding steps. This effect is particularly noticeable on smaller models and in datasets like DevEval. In contrast, AdaDec performs targeted lookahead at every high-uncertainty step, which increases runtime, especially for larger models and more complex DevEval problems, but provides more reliable improvements in functional correctness.
Table 4. Comparison of the average generation time per problem (in seconds) by Greedy, AdapT, AdaDec on HumanEval+, MBPP+, DevEval (the brackets indicate the change ratio relative to Greedy)
Model HumanEval+ MBPP+ DevEval
Greedy AdapT AdaDec Greedy AdapT AdaDec Greedy AdapT AdaDec
DS-1.3B 1.15 2.42 1.46 0.54 0.63 0.97 4.74 9.61 10.24
DS-6.7B 2.37 3.68 3.44 1.02 4.24 1.30 15.54 11.55 26.08
ST-3B 1.70 3.02 2.06 0.59 0.81 0.76 6.87 8.00 16.72
QW-0.6B 6.22 4.07 5.25 10.02 6.14 2.89 20.02 13.86 21.18
QW-1.7B 8.62 4.23 5.27 2.39 2.54 2.28 18.21 13.96 30.33
QW-4B 5.41 7.57 8.06 4.09 3.76 3.37 28.17 17.94 43.97
QW-8B 5.39 4.73 6.94 2.84 4.83 4.01 64.56 16.80 68.08
AVG. $\Delta$ -0.16 +0.23 +0.21 -0.84 -9.48 +8.36
Summary: AdaDec maintains high efficiency by selectively applying pause-and-rerank at a small fraction of high-uncertainty steps, with pause rates mostly below 10%. While this targeted lookahead adds moderate latency compared to Greedy, it remains manageable and effectively balances quality and efficiency. AdapT can sometimes reduce runtime via stochastic sampling, but AdaDec delivers more consistent improvements in functional correctness.
4.4. RQ3: Ablation Study
We perform an ablation study to evaluate the key components of AdaDec. In particular, we investigate whether the learned entropy threshold $\tau^{\bm{LM}}$ serves as an effective signal, both in terms of predictive ability (as a classifier) and in practical decoding performance.
Predictive Effectiveness of Learned Entropy Thresholds. Table 5 summarizes the performance of a logistic regression classifier trained with learned entropy thresholds ($\tau^{\bm{LM}}$) across various language models. Overall, the classifier demonstrates strong effectiveness in predicting whether the top-1 token matches the ground truth. The highest performance is observed on DS-6.7B, achieving an accuracy of 91.05%, an F1 score of 0.9512, and an AUC of 0.9305. Other models, including DS-1.3B and ST-3B, also perform well, with F1 scores above 0.94 and AUC values exceeding 0.89. These results support the effectiveness of learned entropy thresholds as a reliable and lightweight signal for triggering pauses and reranking.
Table 5. Evaluation of logistic regression models trained to predict whether the ground truth token is the top-1 prediction. The learned $\bm{\tau}^{\bm{LM}}$ denotes the entropy threshold selected via logistic regression.
Model Accuracy Precision Recall F1 Score AUC Learned $\bm{\tau}^{\bm{LM}}$
DS-1.3B 0.8994 0.9251 0.9640 0.9442 0.9268 0.9850
DS-6.7B 0.9105 0.9276 0.9761 0.9512 0.9305 1.1338
ST-3B 0.9038 0.9188 0.9780 0.9475 0.9222 1.1005
QW-0.6B 0.8512 0.8774 0.9498 0.9122 0.8914 1.3213
QW-1.7B 0.8664 0.9003 0.9457 0.9225 0.8964 0.6153
QW-4B 0.8807 0.9166 0.9471 0.9316 0.9073 0.7134
QW-8B 0.8843 0.9158 0.9533 0.9342 0.9092 0.7060
AVG. $\Delta$ 0.8852 0.9117 0.9591 0.9348 0.9120 0.9393
Practical Effectiveness of Learned Entropy Thresholds. To further assess the impact of learned thresholds in practice, we compare the decoding performance of models using the learned threshold versus a fixed threshold. The fixed threshold is set to 1.2. Table 6 reports pass@1 results on HumanEval+. The experiments show that using the learned thresholds consistently improves performance compared to the fixed threshold. For DS-1.3B, DS-6.7B, ST-3B, and QW-0.6B, the learned thresholds are close to 1.2, resulting in modest gains. In contrast, for QW-1.7B, QW-4B, and QW-8B, the learned thresholds differ more from 1.2, yielding more noticeable improvements in decoding performance. This confirms that the learned thresholds not only perform well in predictive tasks but also translate into tangible gains during actual decoding.
Table 6. Comparison of Learned vs. Fixed Entropy Thresholds on HumanEval+ (pass@1). Fixed threshold is set to 1.2.
Strategy DS-1.3B DS-6.7B ST-3B QW-0.6B QW-1.7B QW-4B QW-8B AVG
Fixed 57.32% 65.24% 50.00% 21.34% 41.46% 51.22% 61.59% 49.74%
Learned 57.32% 65.24% 50.61% 22.56% 44.51% 56.10% 62.80% 51.31%
$\Delta$ 0.00% 0.00% +0.61% +1.22% +3.05% +4.88% +1.21% +1.57%
Summary: The learned entropy threshold $\tau^{\bm{LM}}$ is both theoretically sound and practically effective: it reliably distinguishes uncertain decoding steps and improves decoding performance when compared to fixed-threshold baselines.
4.5. RQ4: Parameter Sensitivity
To investigate the impact of lookahead length $L$ on decoding performance, we conduct experiments on HumanEval+ using DS-1.3B as the backbone model. We vary $L$ from 2 to 9 and report pass@1 results. Figure 6 illustrates the trend.
Figure 6. Pass@1 Performance of the DS-1.3B Model on the HumanEval+ Dataset Under Different Lookahead Length (L) Configurations.
The results show that pass@1 increases steadily with larger $L$ up to $L=5$, at which point it reaches the peak performance. Beyond $L=5$, the performance declines slightly and then plateaus. This behavior can be explained by the trade-off between local guidance and global noise: shorter lookaheads may fail to capture sufficient future context, while excessively long lookaheads introduce noise and dilute the relevance of local signals. Importantly, across all values of $L$, AdaDec consistently outperforms the baseline decoding strategies, demonstrating the robustness of our approach to the choice of lookahead length.
Summary: The lookahead length $L$ exhibits a non-monotonic effect on performance: moderate values (e.g., $L=5$) maximize the benefits of lookahead guidance, while excessively small or large values reduce effectiveness. Notably, AdaDec consistently surpasses baseline performance regardless of the choice of $L$.
5. Discussion
5.1. Limitations
While AdaDec demonstrates improvements in code generation accuracy and maintains acceptable computational cost, it has several limitations that warrant further investigation:
Reliance Solely on Shannon Entropy. Our uncertainty trigger is based solely on the Shannon entropy of the token distribution, which provides a coarse measure of model uncertainty. More nuanced contextual features—such as entropy of subpopulations of tokens, token-level gradient norms, or higher-order statistics—might better capture subtle divergences. Future work could integrate richer uncertainty signals or leverage learned uncertainty estimators.
Threshold Learning Simplicity. We employ a logistic regression classifier to learn entropy thresholds, which offers interpretability but may be insufficiently expressive for capturing complex interactions among uncertainty features. More powerful models (e.g., small neural networks) could potentially improve threshold accuracy, at the cost of increased complexity.
Potential Bias from Training Data. Although we carefully avoided direct overlap between training and evaluation datasets, our threshold learning still relies on synthetic data generated from a specific corpus. If the distribution of training data diverges significantly from the deployment scenario, the learned thresholds may fail to transfer robustly. Future work could explore continual adaptation or online threshold adjustment.
Addressing these limitations will be an actionable direction for future work, towards making uncertainty-guided adaptive decoding more robust, flexible, and broadly applicable.
5.2. Case Study
To demonstrate how AdaDec enhances decoding decisions in practice, we present a representative example from the HumanEval+ benchmark (Figure 7). The task is to implement the function check_dict_case, which returns True if all dictionary keys are either lowercase or uppercase, and False otherwise. The prompt provides detailed instructions along with multiple illustrative examples.
Figure 7. A case study from HumanEval+
At a certain point, the model DS-1.3B has produced the initial structure shown in the “Current Seq” portion. At this stage, it must decide how to proceed. A standard greedy decoding checks islower() and isupper() directly on all keys, implicitly assuming that the keys are strings. This assumption is unsafe: if any key is a non-string (e.g., an integer), a runtime error will occur. In contrast, AdaDec detects high entropy at this decision point and triggers lookahead reranking. Among the candidate continuations, it selects a branch that first verifies whether all keys are strings before checking their case. This results in a more robust and semantically correct implementation that gracefully handles edge cases, faithfully reflecting the prompt’s intent.
This example illustrates how AdaDec uses entropy signals and semantic reranking to proactively avoid subtle bugs that are often missed by standard decoding strategies. As a result, it improves the reliability of generated code without requiring any changes to the underlying model.
6. Threats to Validity
We analyze threats to validity across four dimensions—internal, external, construct, and conclusion—and note our mitigation strategies or key limitations.
Internal Validity. Drift point identification may be imperfect: our semi-automated AST comparison could miss subtle semantic shifts (e.g., identifier drift), and human verification introduces potential annotator bias. We mitigated these risks through clear guidelines and cross-checking. Entropy-based analysis may also depend on tokenizer or model-specific distributions; to reduce this risk, we validated results across multiple models and confirmed entropy–rank correlations in RQ2.
External Validity. We focus on open-source models (DeepSeek, StableCode, Qwen3) and three benchmarks (HumanEval+, MBPP+, DevEval). Results may not fully generalize to proprietary models (e.g., GPT-4, Codex) or to broader real-world tasks. Future work should extend evaluation to more diverse datasets and model families.
Construct Validity. We evaluate functional correctness (Pass@1) and efficiency (latency), which directly measure the primary objectives of correctness and performance. Nevertheless, these metrics do not capture other important qualities such as readability or adherence to coding conventions. Similarly, our use of Shannon entropy as an uncertainty signal assumes it faithfully reflects model uncertainty, though it may only provide a coarse approximation. Alternative measures (e.g., calibration error, variance across sampling runs) could offer complementary insights.
Conclusion Validity. Conclusion validity threats arise from the statistical robustness of our empirical results. Although we report consistent trends across multiple models and datasets, the sample size of benchmarks such as HumanEval+ remains relatively small, which may limit the statistical power of our findings. Additionally, hyperparameter choices (e.g., lookahead length $L$, candidate pool size $B$) may affect performance outcomes, and while we explored reasonable ranges, further sensitivity analysis would strengthen confidence in our conclusions.
7. Related Work
7.1. LLM-based Code Generation
Large Language Models (LLMs) have shown strong potential in software engineering tasks such as code generation (Li et al., 2023; Di and Zhang, 2025; Tian et al., 2024; Jiang et al., 2025; Lin et al., 2024), unit test generation (Nan et al., 2025; Yang et al., 2024; Yuan et al., 2024; Lops et al., 2025; Ryan et al., 2024), and comment generation (Haider et al., 2024; Katzy et al., 2025; Lu and Liu, 2024). Typically, LLMs generate code from textual requirements via prompt-based interaction. By learning expressive representations from large-scale code and documentation corpora, they can produce syntactically correct and context-aware code snippets (Yuan et al., 2023; Sun et al., 2024; Ugare et al., 2024). Open-foundation models further advance tasks like code completion, bug fixing (Fu et al., 2022; Jin et al., 2023; Xia and Zhang, 2022), and test-case generation.
To evaluate their effectiveness, benchmarks like HumanEval (Chen et al., 2021b), MBPP (Austin et al., 2021), and DS-1000 (Lai et al., 2022) focus on function-level generation, while ClassEval (Du et al., 2023) extends evaluation to class-level tasks. More recently, CoderEval (Zhang et al., 2024) and DevEval (Li et al., 2024) target repository-level scenarios, reflecting growing research interest in applying LLMs to real-world software development.
7.2. Decoding strategies for LLMs
Large Language Models (LLMs) generate text in an autoregressive manner, producing one token at a time. At each generation step, the model samples a token from its predicted probability distribution. The most straightforward sampling strategy is greedy decoding, which selects the token with the highest probability as the next word. A more refined variant is beam search (Freitag and Al-Onaizan, 2017), which maintains multiple candidate sequences at each decoding step and preserves the top-k sequences based on cumulative log-probabilities.
Beyond these, commonly used sampling strategies include temperature sampling (Holtzman et al., 2020), top-p (nucleus) sampling (Holtzman et al., 2020), and top-k sampling (Holtzman et al., 2020), each designed to balance diversity and coherence in generated outputs.
To further improve sampling for code generation tasks, Li et al. introduced AdapT Sampling (Zhu et al., 2023) (Adaptive Temperature Sampling). This method dynamically adjusts the temperature parameter T based on token context. For instance, when the current token initiates a code block, a higher T is used to encourage diversity and creativity. In contrast, for other tokens, a lower T is applied to suppress random noise and preserve syntactic and semantic correctness.
Another relevant approach is UnCert-CoT (Zhu et al., 2025), which enhances code generation through uncertainty-aware Chain-of-Thought (CoT) reasoning, activating multi-path reasoning when token uncertainty surpasses a fixed threshold. While effective, its relies on fixed thresholds and costly CoT reasoning. In contrast, AdaDec learns model-specific uncertainty thresholds for stronger generalization, while its lightweight rerank mechanism cuts computational costs significantly without compromising benchmark performance.
8. Conclusion
This paper presents AdaDec, an uncertainty-guided adaptive decoding framework for large language model-based code generation. Motivated by an empirical study revealing that ranking errors at high-uncertainty decoding steps often lead to logic drift, AdaDec introduces a novel pause-then-rerank mechanism that dynamically identifies and corrects uncertain token predictions using Shannon entropy. By learning model-specific entropy thresholds and employing a lookahead-based reranking strategy, AdaDec achieves substantial improvements in accuracy while maintaining computational cost and efficiency within acceptable bounds. Experiments on HumanEval+, MBPP+, and DevEval benchmarks demonstrate that AdaDec consistently outperforms all compared methods. These findings underscore the importance of token-level uncertainty modeling and open up new directions for adaptive decoding in code generation and beyond.
9. Data Availability
To facilitate the replication study, we have released our data and code at : https://anonymous.4open.science/r/AdaDec-657F.